### v1

# Visual Writing: Writing by Manipulating Visual Representations of Stories

Abstract.
We introduce “visual writing”, an approach to writing stories by manipulating visuals instead of words. Visual writing relies on editable visual representations of time, entities, events, and locations to offer representations more suited to specific editing tasks. We propose a taxonomy for these representations and implement a prototype software supporting the visual writing workflow. The system allows writers to edit the story by alternating between modifying the text and manipulating visual representations to edit entities, actions, locations, and order of events. We evaluate this workflow with eight creative writers and find visual writing can help find specific passages, keep track of story elements, specify edits, and explore story variations in a way that encourages creativity.
1. Introduction
Writing stories and exploring alternative narratives can be challenging. Although using large language models (LLMs) can help writers (Chakrabarty et al., 2024), it remains difficult for users to formulate their intent using words (Dang et al., 2022b) and review the generated text to ensure it aligns with their intent (Masson et al., 2024).
Consider a task as simple as changing the location of a character in a story: the cat no longer goes to the barn but instead wanders about the lake. Using “replace all” instances of ‘barn’ is insufficient because substituting a location comes with implications. For example, descriptions and actions pertaining to the barn would need to be updated to reflect the new context of being by the lake. Further, not all instances necessarily need to be impacted. The fact that the cat no longer visits the barn should not result in the barn being completely removed from the story. Instead, writers must carefully plan to find all and only the references to the location associated with the character’s change of action and make tedious revisions to the text to maintain spatial and semantic consistencies. This busywork is prone to errors, and writers often resort to creating external documents such as maps, spreadsheets, and timelines that they painstakingly update to keep track of the characters, events, and locations (Marshall, 2014; Ackerman, 2017; Neuwirth and Kaufer, 1989; SaveTheCat, 2014). While LLMs could help with this busywork, expressing distinct changes can be challenging with prompts (Masson et al., 2024), and the extent and suitability of the changes made by the generative model could be difficult to verify.
The challenge stems from the mismatch between the reasoning and representation spaces: in our above example, the writer performs spatial reasoning using a textual representation. A spatial representation, such as a map of the story world, would be more suited for such a task, where changing the location of the character becomes as simple as moving its position on the map. In text, edits relevant to space and locations in the story are comparatively challenging to accomplish, which can be a deterrent for writers wanting to change the narrative and explore variations. And this same argument applies to all other aspects of a story such as time, events, and characters, for which we posit superior representations can be designed.
We define visual writing as the use of visual representations of story elements as an input medium for expressing writing intents. For each story element, we explore the use of a visual representation that conceptually aligns with that element. Depending on their editing task, users can decide to edit the text of the story, in which case the visual representations are updated to match the new text. Alternatively, they can manipulate the visual representations through direct manipulation to suggest changes to the text and rapidly explore variations. We discuss a taxonomy of these visual representations, drawing on the story elements framework from narratology (§3). In a prototype software (§4), we implement three of these visual representations: (1) a diagrammatic view of the entities in the story and their relationships, allowing to add, remove, and edit characters and add new actions between them (Figure 1.c); (2) a spatial view of the locations in the story and the locations of the entities, allowing to add new locations and move entities (Figure 1.d); and (3) a timeline of the events in the story, allowing to reorder the events, quickly find specific scenes, and precisely modify the entities and locations of the selected scenes by leveraging the other two views (Figure 1.b).
We validate this approach in a user study with eight creative writers (§5). Results show participants appreciated the visual representations to locate scenes and keep track of story elements while writing. They were successful in expressing their editing intents through direct manipulations of the visual elements. Furthermore, participants appreciated this approach to edit text from an exploration and creativity perspective. This points to the potential of visual representations for writing and calls for more work in that direction.
2. Related Work
Visual writing draws inspiration from work on visualizing stories and efforts to give alternative representations to edit content. Below, we review the literature on these topics.
2.1. Visualizing Stories
Visual representations of stories and story patterns have a long-standing presence in writing, with staple examples including Freytag’s pyramid story (Freytag, 1895) and abstract diagrams capturing the “shape of the story” (Pope, 1998; Comberg, [n. d.]; Reagan et al., 2016). Such visual representations are not only useful for teaching the craft of storytelling, they are also considered as useful aids to support the writing process. Referring to the visual representations, screenplay writer E. Williams says: “I find story shape to be a great tool […] to make sure my story holds together” (Williams, [n. d.]).
There exist many techniques to visualize stories, spanning from classical charts like word clouds, to more artistic visualizations like calligrams (Brath, 2021). Often, these visualizations summarize entire stories in an attempt to highlight a specific facet (e.g., emotions (Maharjan et al., 2018) or topic recurrence (Zhu, 2013)). For example, Storyline visualizations (Munroe, 2009; Tanahashi and Kwan-Liu Ma, 2012) depict the evolution of the relationships between characters through timelines. StoryCake (Qiang and Bingjie, 2016) uses a hierarchical plot to visualize the structure of nonlinear stories. Story Curves (Kim et al., 2018) visualizes non-linear narratives by showing both the order in which the events are told and the order of the underlying story. StoryPrint (Watson et al., 2019) shows the scenes and character emotions through circular timelines. Each of these visual representations presents the ability to support readers, writers, and analysts alike in appreciating or even analyzing the stories from a specific lens.
Other story visualizations have been used in tools to help during script and story writing. For example, CARDINAL (Marti et al., 2018) is a tool to help screenwriters visualize their script through different views, including a visualization of the interactions between characters and an animated 2D and 3D views to play the scene and visualize locations and movements of the characters. Similarly, Portrayal (Hoque et al., 2023) attempts to help writers develop their characters by extracting and presenting information related to the characters in interactive visualizations.
Our work explores the potential of expanding the role of such visualizations to also be manipulable. That is, by mapping visualization elements directly to entities, events, and locations in the text story, we explore using visualizations as the medium to modify the story. This approach is motivated by the potential of visual representations to express intent, which we discuss next.
2.2. Editing Content Through Alternative Representations
Programming might be the domain that most embraced the use of alternative representations. Approaches include visual programming (Myers, 1986; Burnett and McIntyre, 1995), which allows generating programs by manipulating elements in more than one dimension, code projections (Gobert and Beaudouin-Lafon, 2023; Ko and Myers, 2006) which augment programming languages with tools and widgets that are alternative representations of fragments of code, and bi-directional programming (Mayer et al., 2018; Hempel et al., 2019) where the output of a program can be manipulated to edit the program itself. Similar to our work, the motivation behind these approaches is that some tasks are easier done using different representations: it is easier to find a colour using a colour picker than by typing its hexadecimal code, and it is easier to move an object on the screen through direct manipulation (Shneiderman, 1983) than by typing its coordinates.
In the creative writing domain, this idea of manipulating projections or visual representations to edit stories is largely untapped, which is surprising given the reliance on visual representations to capture story concepts (Freytag, 1895; Comberg, [n. d.]). Instead, most writing support tools ask users to manipulate text using text, typically through prompts or suggestions (Calderwood et al., 2020; Buschek et al., 2021; Yuan et al., 2022; Dang et al., 2023; Lee et al., 2024; Reza et al., 2024). There are some exceptions, such as the proposal of Kempen et al. (1987) to represent and manipulate text in different ways, such as manipulating the syntax tree to allow restructuring sentences. Similarly, Arnold et al. (2021) suggested that words could be directly manipulated to be reordered. Beyond syntactic level modifications, Dang et al. (2022a) proposed a system that allows reorganizing, merging, and removing paragraphs by manipulating textual summaries instead of the text. In the same vein, VISAR (Zhang et al., 2023) automatically generates an interactive visual outline of an argumentative text, which can be edited to change the structure of the argument. Finally, TaleBrush (Chung et al., 2022) leverages a canvas where sketching allows modulating the fortune of characters in a story.
Our work contributes to this line of research by formalizing the concept of visual writing. The concept is rooted in the same principles as visual programming and code projections, positing that elements of a story become easier to understand and manipulate when presented in alternative forms. Specifically, we design visualizations grounded in narratology to explore this approach.
3. Taxonomy of Visual Representations
Narratology separates the chronological sequence of the story (fabula) from the narrative sequence (syuzhet) of the story. At its core, a fabula consists of four main story elements: events, actors111In narratology, actors refer to “agents that perform actions. They are not necessarily human” (Bal, 1997). To avoid confusion, we use the term entity instead throughout this paper., time, and location (Bal, 1997). Similarly, we define visual writing as the use of visuals that represent and manipulate these four elements.
In their simplest form, visual representations can depict elements one at a time. For example, a list of locations would allow adding, removing, and editing the characteristics of each location. And a list of entities, similar to the character sheets writers often create to plan their stories (SaveTheCat, 2014), would allow adding and removing entities. However, many important aspects of stories are encoded through the interaction between elements (e.g., the time of an event or the location of an entity). Below, using the narratology lens, we discuss a non-exhaustive taxonomy of possible visual representations as an initial guide to inform the design of tools supporting visual writing.
Spatial representations depict elements based on their location. For example, a spatial representation could be a map with dots, where each dot represents where events happened and where entities are located. Or it could be more abstract and only represent distances between elements. Spatial representations are used by systems like CARDINAL (Marti et al., 2018) to show the relative positioning of characters within a scene.
Event-driven representations depict elements based on the events they relate to. For example, a node graph could be an event-driven representation if each node represents an event and the edges encode another element such as the entities (in which case, connected nodes would represent events affecting the same entity) or time (in which case the directed edges would indicate the flow of the events). Some examples of event-driven representations are lists of outline points and event graphs.
Entity-driven representations depict elements based on the entities they relate to. For example, a node graph could be an entity-driven representation if each node represents an entity and the edges encode another element such as the events (in which case, each edge is an event connecting the entities involved in the event) or time (in which case, the directed edges indicate the order in which the entities intervene in the story). Entity-driven representations, such as character sheets and family trees, are often created by writers to keep track of character details (SaveTheCat, 2014).
Temporal representations depict elements based on the time they happen in the fictional world of the story or in the order they are presented in the text. For example, a timeline could arrange elements such as events (in which case, the events would be positioned based on when they occur) or entities (in which case, entities would be positioned based on when they appear or when they perform an action). Temporal representations have been more commonly used by previous work on visualizing stories such as Storylines (Munroe, 2009), StoryCake (Qiang and Bingjie, 2016), and StoryPrint (Watson et al., 2019).
The design space for the above categories is vast, and there exist many visualizations to bear inspiration from. Our goal is not to argue which specific design is best—we leave it to future work to explore different variations within these categories (and beyond). Instead, our work focuses on the potential of visual writing, i.e. the use of alternative representation of stories, as an input medium for expressing intent.
4. The Visual Writing Prototype
To test the potential of visual writing, we developed a prototype system, shown in Figure 1, implementing three visual representations: a diagrammatic view of the entities and actions (Entity–Event), a spatial view of the locations and entities (Location–Entity), and a timeline of the events (Time–Event). The focus of the prototype was to help writers explore variations and perform spatial and temporal edits via visual writing. As such, the system allowed specifying edits through direct manipulations of the graphical elements (Shneiderman, 1982; Hutchins et al., 1985) and followed the principles to design creativity support tools (Shneiderman et al., 2006). Below, we detail the design and features of the system. All the examples and screenshots were taken using the system on the first three paragraphs of the story Alice’s Adventures in Wonderland by Lewis Carroll.
4.1. Entities and Actions View
This event-driven representation shows all the entities in the story as nodes with their name and an emoji representing them (Figure 1.c). The actions between the entities are represented as directed edges with a label describing the action in one or two words. By default, the nodes are arranged using a force layout to prevent overlaps, but users can rearrange the position of the entities by dragging them.
4.1.1. Editing the Traits of an Entity
Selecting an entity expands its node to reveal traits related to the entity similar to a list of personality traits which would be present on a character sheet (e.g., “curious”, “adventurous”, “chatty”). The intensity of each trait is rated from 1 to 10. This intensity can be changed by moving the slider beside the trait. This results in editing the text of the story to reflect the new intensity of the trait (Figure 2.a).
4.1.2. Adding and Removing an Entity
Adding an entity is done by double-clicking anywhere on the canvas which opens a text input box to indicate the name of the new entity. Once the name is entered, the entity is created at the location of the click. Removing an entity is done by selecting the entity and then pressing “delete” or “backspace” on the keyboard. This results in updating the text of the story so that the entity is effectively removed (Figure 2.b).
4.1.3. Adding and Removing Actions between Entities
Adding an action is done by creating an edge between two nodes. The direction of the edge indicates the source and target of the action. An action can also be created for the same character by connecting a node to itself. This helps represent actions done by an entity with no target (e.g., walking or thinking). An action can be edited at all times by double-clicking its label and changing its description. Similarly, an action can be removed by being selected and then pressing “delete” or “backspace”. As soon as the edge is created, edited, or deleted, the story text is updated to reflect the change (Figure 2.c).
4.2. Locations and Entities View
The location view is a spatial representation accessed through a tab (Figure 1.d). It shows all the locations in the story as nodes with their name and an emoji representing them. Similar to the entities and actions view (Section 4.1), the view shows all entities as smaller nodes, but their position is conditioned by their location in the story. Entities that go to different locations in the story are duplicated to be represented in all of their locations.
4.2.1. Moving an Entity
Entities can be dragged and moved around the view. If the entity is released on top of a location, then that entity is moved to that location (Figure 3.b-c). Otherwise, the entity goes back to its original location. Once the entity is moved, the story text is updated to reflect the change in location.
4.2.2. Adding a Location
A new location can be added by double-clicking anywhere on the canvas (Figure 3.a). This opens a text input field to specify the name of the location. Once the location is created, entities can be moved to it.
4.3. Timeline of Events View
The timeline view is a temporal representation showing the events in the story represented as vertical lines with the emojis of the entities involved in the event on either side of the line (Figure 1.b). These lines are organized one after the other based on the order in which they are presented in the text. The timeline view is always visible. It allows selecting one or multiple events, in which case the other views are updated to fade out the entities and locations not involved in the selected events. Selecting events also forces future modifications to impact only the selected events.
4.3.1. Reordering Events
Events can be dragged to be reorganized temporally. When an event is dragged, it slides horizontally and can be placed between other events (Figure 4.b-c). Once the event is moved, the story text is updated so that the event happens when indicated in the timeline.
4.3.2. Selecting Events
The timeline also allows finding specific scenes and selecting them to allow precise editing. Hovering over events highlights the corresponding sentence in the story (Figure 4.a) and the corresponding entities and locations in the event and location views. Clicking selects the event, and dragging selects multiple events. Once one or multiple events are selected, any visual edits made in the views will target these selected events. For example, selecting an event and then moving an entity in the location view will update the story text so that the entity moved exactly during the selected event.
4.4. Bi-directional Editor
The rest of the interface is composed of a text editor on the left, a history tree at the bottom to navigate different versions of the story, and interface buttons at the centre to refresh the visuals or rewrite the story (Figure 1.a).
4.4.1. Highlighting Visual Elements and Sentences on Hover
Similar to how hovering over the events in the timeline highlights the sentences, hovering over the sentences in the text editor highlights the corresponding events, entities, and locations in the different views. This allows for quickly finding different scenes. Similarly, placing the caret in one sentence will act as if the events in that sentence were selected in the timeline. As such, future edits to the visual representations will be restricted to that sentence.
4.4.2. Updating the Visual Representations from the Text
When the text is updated manually, the visual representations might become desynchronized, showing the story as it was before the modification. In this case, the button to refresh the visuals becomes blue to indicate that the visual model may need to be updated. Clicking it results in re-extracting the information from the text and refreshing the views.
4.4.3. Writing the Story from the Visual Representations
By clicking the “refresh from visuals” button, the story is rewritten from scratch using only the visual representations as reference. As such, the story preserves the same events, locations, and entities, but everything else is rewritten. This allows specifying the skeleton of a story and quickly exploring different ways of phrasing the story.
4.4.4. Track changes
By default, all changes made by editing the visual representations are tracked, such that removed text is struck in red while additions are shown in blue. This allows writers to easily locate and review the changes before accepting (or rejecting) them.
4.4.5. History Tree
To make edits easily reversible and facilitate the exploration of alternatives, the system also implements a history tree capable of storing different versions of the story and returning to them at any time. Every time a modification happens, be it by editing the visual representation or editing the text of the story, a snapshot of the story and visual representation are added to the tree. If the writer selects a previous snapshot, then the story is changed to that snapshot. If, after selecting a previous snapshot, a modification is made to the story, then this creates a new branch to the story. Then, writers can navigate between the branches and different versions by clicking on the nodes of the history tree.
4.5. Implementation
We implemented the visual writing system using TypeScript and React (React, 2013) for the interface, NextUI (NextUI, 2021) for the graphical components, Slate.js (SlateJS, 2016) for the text editor, and the OpenAI library for prompting large language models (OpenAI, 2023). Under the hood, modifications to the visual representations are turned into engineered prompts sent to OpenAI’s “GPT4-o” model.
4.5.1. Extracting Entities, Locations, and Events
The visual writing approach requires a robust, exhaustive, and fast extraction of the information needed for the visual representations. In our tests, a naive approach consisting of giving a large language model the whole story and asking to extract the necessary information failed at all three of these requirements. Instead, we force the model to output JSON with the proper schema to make the extraction robust. We check that the schema is correct and prompt the model again if necessary. To make the extraction exhaustive, we proceed in multiple steps by first asking the model to extract the entities and locations in the story. We then ask the model to extract events between these entities and at those locations. Finally, and most importantly, we split the whole story into sentences to make the extraction fast. We then prompt the model with the story, but we ask it to extract only the information within the sentence. This allows us to execute the prompt requests in parallel. As such, a text with 30 sentences will send 30 different smaller prompts to the model. This also gives us an indication of the position of the extracted information, as we know that it comes from the sentence being extracted.
4.5.2. Engineered Prompts to Edit the Story
The system relies on many engineered prompts such as splitting requests to make them faster, giving the model only a subset of the text to force localized edits, and giving the model the previous abstract representation of the story and the new one to help it understand the edits. We encourage interested readers to review the source code for more details.
5. User Study
We evaluated the visual writing workflow by conducting a user study using our prototype system. Specifically, the study tried to answer the following questions: Are the visual representations understood and helpful? Can people express their editing intents using visual writing? What is the impact of visual writing on exploration and creativity?
Our study is separated into two parts. The first part isolates the use of the three different views in a controlled setting. Because our approach is compatible with other ways to edit text (e.g., no assistance or LLMs) and our main interest is to measure the usefulness of visual writing, we decided against comparing our approach to a baseline. Instead, we collected qualitative insights through semi-structured interviews and explicitly asked participants to compare visual writing with their own workflow. Representations were also evaluated quantitatively using simple usability questionnaires.
The second part of our study was a free-form creative writing task to let participants appropriate the tool and its features.
5.1. Participants
We recruited 8 creative writers from our social network and word of mouth (age range: 20 to 31, M=24; 7 self-identified as female and 1 as male). All had several years of experience with creative writing, including personal writing (N=7), fiction (N=4), scriptwriting (N=2), and poetry (N=2). On a scale from 1-never to 5-always, they reported how often they use visuals (e.g., mind maps, timelines) during their writing process (Mdn=2), how often they find visuals to help during the writing process (Mdn=3), and how frequently they use AI services to help during their writing activities (Mdn=1). Participants reported using a range of types of visuals to help their writing, including mood boards (N=4), timelines (N=3), mind maps (N=2), and reference imagery (N=2).
5.2. Apparatus
Participants joined the study remotely via the Zoom conference software. They accessed the experiment website using their own computer. The sessions were audio and screen recorded, and interactions with the experiment website were also recorded (e.g., features used, commands executed, and answers to questions). The sessions were about 45 minutes long. In appreciation for their time, participants received a CAD$15 gift card. The study was approved by our institutional ethics board.
5.3. Design and Procedure
Introduction (~2min)
After being greeted, the participant was asked to complete the consent form and demographics questionnaire. They were then told that they would use a system to help them edit stories.
Video Tutorial ((1 + 3 tasks) 30 seconds)
Before each task, the participant watched a short video demonstrating the features they were about to use in the task. The video used a different text and different tasks to make sure the participant could not just reproduce what happened in the tutorial. Before the first task, the participant watched an additional video presenting the concept of the tool and the organization of the interface.
Story Editing Tasks (3 tasks 5min)
Immediately after watching the video tutorial, the participant were prompted to complete the task. We had three task blocks for each of the visual representation views: (1) entities and actions view tasks required removing an entity and then adding a new entity and an action with the main character; (2) location and entities view tasks required moving an entity to an existing location and then creating a new location and move another entity to it; (3) timeline view tasks required the reordering of events. The instructions for each task were phrased as questions and in a way that forced participants to engage with the story. For example, one of the tasks related to the timeline read “What if Jack loses his hat when he flies into the sky?”, forcing participants to understand enough about the story to know where those events occur and move the action to the right location. The tasks were counter-balanced across participants.
Post-Task Questionnaire (3 tasks 2min)
After each task, the participant first indicated how successful they felt in accomplishing the task on a 5-point semantic differential scale. Then, they answered five questions on a 5-point scale related to usability, understandability, and integration into their workflow.
Free-form Exploration (5-15min)
After accomplishing all the tasks and having a good grasp of the tool, the participant engaged in a free-form exploration. We provided the beginning of a story and asked the participant to continue the story with no instructions other than to “explore what could happen next”.
Post-Study Questionnaire (2min)
After completing the free-form part of the study, the participant answered one last questionnaire, which included the creativity support index (CSI (Cherry and Latulipe, 2014)) and the raw NASA TLX.
Semi-structured Interview (10min)
At the end, the participant engaged in a semi-structured interview with the experimenter. After giving their immediate impressions of the tool, they were asked about their experience with respect to the main categories of the CSI, like enjoyment, exploration, expressiveness, immersion, and satisfaction given the effort. They were also asked about each representation and their perceived overall advantages and disadvantages of the system.
5.4. Results
Participants completed the task and the free-form exploration without major difficulties. Below, we summarize the quantitative findings using descriptive statistics. Following previous recommendations for study designs similar to ours (Zhu and Kolassa, 2018; Masson et al., 2023), confidence intervals are calculated using the studentized bootstrapping method using 10,000 repetitions.
Success rate
On a 5-point semantic differential scale from 1-unsuccessful to 5-successful, participants rated their success M=4.25 (95% CI: [3.8, 4.6]) at accomplishing the tasks. Specifically, they were most successful at the task about adding and removing entities (M=4.5, 95% CI: [3.8, 5]), followed by changing the locations of entities (M=4.25, 95% CI: [3.4, 4.8]), and reordering events (M=4, 95% CI: [2.8, 4.7]).
Creativity Support Index (CSI)
Participants generated an average CSI score of 71.5 (SD=21.12) for the system and the tasks. Table 1 shows the breakdown of the CSI results. Enjoyment was both the most highly rated factor and also the factor considered most important by participants. Immersion was ranked second most important yet received the lowest score. Possible reasons for these ratings are discussed in Section 6.
| Scale | Avg. Score (SD) | Avg. Counts (SD) |
| Enjoyment | 15.73 (1.84) | 3.50 (1.00) |
| Immersion | 12.18 (3.04) | 3.13 (1.76) |
| Results Worth Effort | 15.28 (2.26) | 3.00 (1.50) |
| Exploration | 13.48 (2.91) | 2.88 (1.27) |
| Expressiveness | 13.70 (2.86) | 2.50 (0.50) |
| Overall CSI Score | 71.50 (21.12) |
Workload (Raw NASA-TLX)
Participants reported high performance (M=7.49 SD=1.95) for low mental demand (M=4.11, SD=2.38), effort (M=3.39, SD=2.17), frustration (M=2.5, SD=1.2), temporal demand (M=1.18, SD=1.04), and physical demand (M=1.18, SD=1.04).
Questionnaire
Overall, answers to the 5-point scale statements from 1-strongly disagree to 5-strongly agree were mostly positive, see Figure 5 for a breakdown of the results. Participants found the timeline the easiest view to understand (Mdn=4) and use (Mdn=4). Participants mostly agreed that the views helped accomplish the tasks (Mdn=5), that they would be useful within their writing process (Mdn=4), and that they would easily be integrated into their writing process (Mdn=4).
6. Discussion
We first discuss our results in light of participants’ comments before opening the discussion to other themes stemming from our observations and discussions with participants.
Visual writing can help find scenes and keep track of events, entities, and locations
Most participants agreed that “The visual representation was easy to understand” for all three views. Particularly, they mentioned how the views helped them. P5 appreciated that you can “click on a person, click on a location, and then see stuff pop up” instead of “having to look through […] many paragraphs […] or doing keyword search”. P4 explained “Just having a visual representation helps, you know, just with the different perspectives”. P3 added “Moving around the timeline was a good way of, kind of figuring out different events”. P1 and P5 also mentioned how they struggle to keep track of everything while writing their stories and that the system would help them keep track of that. For instance, P5 details that currently they “have booklets everywhere, and [they] have it annotated weird […] it’s a mess and this is so much easier. You got like, you know, your locations, your people […] keeping track of things. Organizing who’s where, what’s going on, being able to scroll through, like highlighting part of the text shown on the timeline […], making it easy to understand”.
Visual writing can help specify temporal, spatial, and entity-related edits
Participants reported high success rates when accomplishing the tasks and provided high ratings for statements such as “The visual representation helped me accomplish the tasks” and “The visual representation was easy to use”. This suggests that visual writing facilitates temporal, spatial, and entity-related edits through the views provided in our system. Participant’s comments further support these results. For example, P2 mentioned “If I add in an extra character, it gives you just enough text to build that story with that new character”. P4 appreciated that it made the story consistent when editing locations: “When I was dealing with the change where the hat was supposed to be left at home, the fact that it also updated the sentence where the hat flew off his head that is extremely useful because that’s the sort of inconsistencies that can easily creep in when you’re making modifications like that”. Similarly, P6 explained, “I really liked the ability to create relations between characters […] It’s something that made it a lot easier than any sort of other systems that I’ve used” and later added “[the tool] could be useful as a touch-up tool for when I want to make larger sweeping changes to a very long existing story […] I can confirm whether or not I want to keep each of these edits.”. This idea of reviewing the changes was important for participants, as P3 explained “You’d still have to review it, but it made enough changes that it [the story] was able to flow well.”
Visual writing can help explore and be creative
Many participants felt the tool was useful as a creative assistant. P6 explained “A lot of times, I am interested in how making a change to a given detail in a story would look […] I was appreciating it [the tool] as sort of an exploratory tool where even if it doesn’t end up producing something that I would personally publish, I think it’s really good to give an example that I can then engage with”. Similarly, P5 mentioned “You can just like, throw in a verb or the general action, interaction between them and it was kind of like spitballing ideas”. Regarding the timeline, P7 felt that “it’s an easy way to explore your options of what it sounds like if something happens here or if it sounds better when it’s there.”. They also mentioned how just looking at the visualization could help, for instance: “In terms of ideas, I could see how looking at the map, something could pop up into your head […] once you have, like, those objects laid out on a map, then that can help you spark inspiration”.
6.1. Observations and Participants’ Comments
6.1.1. Participants had different preferences
For instance, P4 mentioned “I really really like having a visual representation of the story, but this is not the visual representation I would choose […] I would like a combination of those two things [timeline and entity view] where those interlinks and connections appear within the timeline.”. Similarly, P6 explained how a single timeline would not be enough for them: “my brain interprets timelines in stories where I enjoy having things happening in multiple places at the same time and having a single linear approach makes it very hard to have that sort of interesting kind of development”. Other than visual representations, P4 also mentioned they “don’t want the system overwriting […] because then it’s not me expressing myself through text” and instead, they would prefer “just highlighting the places where things need to change, or the things are inconsistent”. This corroborates previous research on creative support tools needing to offer “many paths and many styles” to do tasks because users are unpredictable and have different preferences (Shneiderman et al., 2006).
6.1.2. Participants would have liked more control in using the visual representations to edit the story
While participants could express their intents with the tool, there are nuances in how an idea can be expressed in terms of writing style and words used, which is not currently supported by our system. For example, when discussing adding actions by connecting two entities, P8 mentioned “I feel like that’s very limiting how you wanted to express what you want versus just typing it out in a sentence because it kind of format it in a specific manner”. Similarly, P7 explained “in terms of expressing myself, I would say that it does help […] but I don’t think it allows for like total freedom of expression […] if you had something like, in your mind of how you wanted to say it, it might not come out that way”.
6.2. Limitations
6.2.1. The participants in our study were experienced and hesitant to using AI
All the participants in our study reported having multiple years of creative writing experience. Similarly, on a scale from 1-never to 5-always, participants rated their frequency of use of AI services a median of one, with five participants reporting never using AI services. P4 explained “I write because I want to, so if I want to be writing then there’s ideas that I have, and I want to express them. I don’t want a system to tell me what it is that I am going to be expressing”. This confirms previous findings that writers have different opinions towards AI (Hertzmann, 2018; Biermann et al., 2022) and those opinions might have impacted our results.
6.2.2. The system and the representations do not scale to long stories
One main limitation of the representations used in the system is that they quickly become too crowded and unusable for long stories. Currently, the system solves this problem through filtering mechanisms when the cursor hovers over a sentence or a scene. This results in hiding the elements irrelevant to the passage currently being selected. But this only works for local edits. When the edits need to happen between scenes far apart in the story, then it becomes challenging to find the right scene or the correct location. We believe that, for these cases, new aggregated representations and semantic zooming would be necessary.
6.2.3. The system might not have been robust enough to test the full potential of visual writing
While the system was optimized to be robust, it still depends on a large language model to understand the story and the instructions properly. In our tests, but also during the study, we observed the system sometimes responding unexpectedly by modifying elements it should not or by refusing to do certain modifications. Additionally, despite our best efforts to parallelize requests and speed up the execution, some manipulations were still slow to execute (up to 10-15 seconds at times). These issues might have confounded the results of our study. For example, it is known that to promote user exploration, interactions should be fast and with effects immediately visible (Masson et al., 2022; Shneiderman et al., 2006; Shneiderman, 1983).
6.3. Future Work
6.3.1. Visual writing to edit writing style and story plots
Creative writing is more than events, actors, locations, and time. We believe that visual writing could also help editing other aspects, such as the style of the writing. As proposed by Kempen et al. (1987), the syntax tree of the sentences could be made visible and be manipulable to restructure sentences. Alternatively, the tone, the length of the sentences, or the sentence structure could be represented using histograms or by embellishing the text. Then, manipulating these visuals would allow tweaking the style of the story. At a higher level, visual writing could also help revise the plot of the story. Similar to work on manipulating outlines (Dang et al., 2022a; Zhang et al., 2023), the idea would be to represent the main plot events of the story and then make them manipulable to reorder or change them.
6.3.2. Exploring new visual interactions
The idea of visual writing is broader than the representations and interactions proposed by our system. Any visual representation better suited for a task could be beneficial. For example, one could imagine using images to give the system an indication of what an object, character, or location look like. Then, the image could be modified to update the story (e.g., painting the hair of a character blond would edit the story so that the description matches the new hair colour). The system could also allow highlighting part of the image to indicate which aspect should be more or less emphasized in the description. Similarly, while our implementation uses LLM, visual writing could be implemented without and made more or less autonomous (Morris et al., 2024). Using LLM comes about with many challenges related to biases, privacy, and plagiarism (Weidinger et al., 2021; Meyer et al., 2023). An alternative, suggested by P4, could be to have the visual writing system only highlight the parts that should be modified to help the writer identify and edit them themselves.
6.3.3. Suggestions and creativity support in the visual space
The mental processes involved in interpreting visuals differ from those involved in processing language (Paivio, 1990). Future work could explore the impact on creativity of writing stories visually rather than with text. Similarly, traditional creativity support techniques for writers would be interesting to adapt to the visual space. For instance, there are many benefits of phrase suggestions (Buschek et al., 2021; Bhat et al., 2023), and future work could investigate if there are similar benefits to suggestions in the visual space. These visual suggestions would be shown when an interaction with a visual element is initiated to highlight different alternatives and to “auto-complete” the interaction.
7. Conclusion
We proposed visual writing, an approach to writing by manipulating visual representations to express editing intents. Through a taxonomy of possible visual representations based on narratology and a prototype implementation of some of these visual representations, we demonstrated one possible design for visual writing tools. A study with eight creative writers showed the potential of creative writing to help keep track of story elements, rapidly specify edits, and explore story variations in a way that encourages creativity. Broadly, our work advocates for a new generation of writing support tools that allow text editing through direct manipulation of visual representations.
Acknowledgements.
This work was supported by NSERC under grant RGPIN-2018-05072.References
- (1)
- Ackerman (2017) Angela Ackerman. 2017. The Efficient Writer: Using Timelines to Organize Story Details.
- Arnold et al. (2021) Kenneth C Arnold, April M Volzer, and Noah G Madrid. 2021. Generative Models Can Help Writers without Writing for Them. In IUI Workshops.
- Bal (1997) Mieke Bal. 1997. Narratology: Introduction to the Theory of Narrative (2nd edition ed.). University of Toronto Press, Scholarly Publishing Division, Toronto.
- Bhat et al. (2023) Advait Bhat, Saaket Agashe, Parth Oberoi, Niharika Mohile, Ravi Jangir, and Anirudha Joshi. 2023. Interacting with Next-Phrase Suggestions: How Suggestion Systems Aid and Influence the Cognitive Processes of Writing. In Proceedings of the 28th International Conference on Intelligent User Interfaces (IUI ’23). Association for Computing Machinery, New York, NY, USA, 436–452. https://doi.org/10.1145/3581641.3584060
- Biermann et al. (2022) Oloff C. Biermann, Ning F. Ma, and Dongwook Yoon. 2022. From Tool to Companion: Storywriters Want AI Writers to Respect Their Personal Values and Writing Strategies. In Designing Interactive Systems Conference. ACM, Virtual Event Australia, 1209–1227. https://doi.org/10.1145/3532106.3533506
- Brath (2021) Richard Brath. 2021. Surveying Wonderland for Many More Literature Visualization Techniques. https://arxiv.org/abs/2110.08584v1.
- Burnett and McIntyre (1995) Margaret M. Burnett and David W. McIntyre. 1995. Visual Programming. Computer 28, 03 (March 1995), 14–16. https://doi.org/10.1109/MC.1995.10027
- Buschek et al. (2021) Daniel Buschek, Martin Zürn, and Malin Eiband. 2021. The Impact of Multiple Parallel Phrase Suggestions on Email Input and Composition Behaviour of Native and Non-Native English Writers. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3411764.3445372
- Calderwood et al. (2020) Alex Calderwood, Vivian Qiu, Katy Ilonka Gero, and Lydia B. Chilton. 2020. How Novelists Use Generative Language Models: An Exploratory User Study.. In HAI-GEN+ User2agent@ IUI.
- Carroll et al. (2009) Erin A. Carroll, Celine Latulipe, Richard Fung, and Michael Terry. 2009. Creativity factor evaluation: towards a standardized survey metric for creativity support. In Proceedings of the Seventh ACM Conference on Creativity and Cognition (Berkeley, California, USA) (C&C ’09). Association for Computing Machinery, New York, NY, USA, 127–136. https://doi.org/10.1145/1640233.1640255
- Chakrabarty et al. (2024) Tuhin Chakrabarty, Vishakh Padmakumar, Faeze Brahman, and Smaranda Muresan. 2024. Creativity Support in the Age of Large Language Models: An Empirical Study Involving Emerging Writers. arXiv:2309.12570 [cs]
- Cherry and Latulipe (2014) Erin Cherry and Celine Latulipe. 2014. Quantifying the Creativity Support of Digital Tools through the Creativity Support Index. ACM Trans. Comput.-Hum. Interact. 21, 4 (June 2014), 21:1–21:25. https://doi.org/10.1145/2617588
- Chung et al. (2022) John Joon Young Chung, Wooseok Kim, Kang Min Yoo, Hwaran Lee, Eytan Adar, and Minsuk Chang. 2022. TaleBrush: Sketching Stories with Generative Pretrained Language Models. In CHI Conference on Human Factors in Computing Systems. ACM, New Orleans LA USA, 1–19. https://doi.org/10.1145/3491102.3501819
- Comberg ([n. d.]) David Comberg. [n. d.]. Kurt Vonnegut on Shape of Stories. https://www.youtube.com/watch?v=oP3c1h8v2ZQ
- Dang et al. (2022a) Hai Dang, Karim Benharrak, Florian Lehmann, and Daniel Buschek. 2022a. Beyond Text Generation: Supporting Writers with Continuous Automatic Text Summaries. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology. ACM, Bend OR USA, 1–13. https://doi.org/10.1145/3526113.3545672
- Dang et al. (2023) Hai Dang, Sven Goller, Florian Lehmann, and Daniel Buschek. 2023. Choice Over Control: How Users Write with Large Language Models Using Diegetic and Non-Diegetic Prompting. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–17. https://doi.org/10.1145/3544548.3580969 arXiv:2303.03199 [cs]
- Dang et al. (2022b) Hai Dang, Lukas Mecke, Florian Lehmann, Sven Goller, and Daniel Buschek. 2022b. How to Prompt? Opportunities and Challenges of Zero- and Few-Shot Learning for Human-AI Interaction in Creative Applications of Generative Models. https://arxiv.org/abs/2209.01390v1.
- Freytag (1895) Gustav Freytag. 1895. Technique of the drama: An exposition of dramatic composition and art. S. Griggs.
- Gobert and Beaudouin-Lafon (2023) Camille Gobert and Michel Beaudouin-Lafon. 2023. Lorgnette: Creating Malleable Code Projections. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3586183.3606817
- Hempel et al. (2019) Brian Hempel, Justin Lubin, and Ravi Chugh. 2019. Sketch-n-Sketch: Output-Directed Programming for SVG. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (UIST ’19). Association for Computing Machinery, New York, NY, USA, 281–292. https://doi.org/10.1145/3332165.3347925
- Hertzmann (2018) Aaron Hertzmann. 2018. Can Computers Create Art? https://arxiv.org/abs/1801.04486v6.
- Hoque et al. (2023) Md Naimul Hoque, Bhavya Ghai, Kari Kraus, and Niklas Elmqvist. 2023. Portrayal: Leveraging NLP and Visualization for Analyzing Fictional Characters. In Proceedings of the 2023 ACM Designing Interactive Systems Conference. ACM, Pittsburgh PA USA, 74–94. https://doi.org/10.1145/3563657.3596000
- Hutchins et al. (1985) Edwin L. Hutchins, James D. Hollan, and Donald A. Norman. 1985. Direct Manipulation Interfaces. Human–Computer Interaction 1, 4 (Dec. 1985), 311–338. https://doi.org/10.1207/s15327051hci0104_2
- Kempen et al. (1987) Gerard Kempen, G. Ambeek, Peter Desain, L. Konst, and Koenraad De Smedt. 1987. Author Environments: Fifth Generation Text Processors. (Jan. 1987).
- Kim et al. (2018) Nam Wook Kim, Benjamin Bach, Hyejin Im, Sasha Schriber, Markus Gross, and Hanspeter Pfister. 2018. Visualizing Nonlinear Narratives with Story Curves. IEEE Transactions on Visualization and Computer Graphics 24, 1 (Jan. 2018), 595–604. https://doi.org/10.1109/TVCG.2017.2744118
- Ko and Myers (2006) Amy J. Ko and Brad A. Myers. 2006. Barista: An Implementation Framework for Enabling New Tools, Interaction Techniques and Views in Code Editors. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’06). Association for Computing Machinery, New York, NY, USA, 387–396. https://doi.org/10.1145/1124772.1124831
- Lee et al. (2024) Mina Lee, Katy Ilonka Gero, John Joon Young Chung, Simon Buckingham Shum, Vipul Raheja, Hua Shen, Subhashini Venugopalan, Thiemo Wambsganss, David Zhou, Emad A. Alghamdi, Tal August, Avinash Bhat, Madiha Zahrah Choksi, Senjuti Dutta, Jin L. C. Guo, Md Naimul Hoque, Yewon Kim, Simon Knight, Seyed Parsa Neshaei, Agnia Sergeyuk, Antonette Shibani, Disha Shrivastava, Lila Shroff, Jessi Stark, Sarah Sterman, Sitong Wang, Antoine Bosselut, Daniel Buschek, Joseph Chee Chang, Sherol Chen, Max Kreminski, Joonsuk Park, Roy Pea, Eugenia H. Rho, Shannon Zejiang Shen, and Pao Siangliulue. 2024. A Design Space for Intelligent and Interactive Writing Assistants. https://doi.org/10.1145/3613904.3642697 arXiv:2403.14117 [cs]
- Maharjan et al. (2018) Suraj Maharjan, Sudipta Kar, Manuel Montes, Fabio A. González, and Thamar Solorio. 2018. Letting Emotions Flow: Success Prediction by Modeling the Flow of Emotions in Books. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers), Marilyn Walker, Heng Ji, and Amanda Stent (Eds.). Association for Computational Linguistics, New Orleans, Louisiana, 259–265. https://doi.org/10.18653/v1/N18-2042
- Marshall (2014) Colin Marshall. 2014. How J.K. Rowling Plotted Harry Potter with a Hand-Drawn Spreadsheet | Open Culture.
- Marti et al. (2018) Marcel Marti, Jodok Vieli, Wojciech Witoń, Rushit Sanghrajka, Daniel Inversini, Diana Wotruba, Isabel Simo, Sasha Schriber, Mubbasir Kapadia, and Markus Gross. 2018. CARDINAL: Computer Assisted Authoring of Movie Scripts. In Proceedings of the 23rd International Conference on Intelligent User Interfaces (IUI ’18). Association for Computing Machinery, New York, NY, USA, 509–519. https://doi.org/10.1145/3172944.3172972
- Masson et al. (2023) Damien Masson, Sylvain Malacria, Géry Casiez, and Daniel Vogel. 2023. Statslator: Interactive Translation of NHST and Estimation Statistics Reporting Styles in Scientific Documents. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3586183.3606762
- Masson et al. (2024) Damien Masson, Sylvain Malacria, Géry Casiez, and Daniel Vogel. 2024. DirectGPT: A Direct Manipulation Interface to Interact with Large Language Models. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3613904.3642462
- Masson et al. (2022) Damien Masson, Jo Vermeulen, George Fitzmaurice, and Justin Matejka. 2022. Supercharging Trial-and-Error for Learning Complex Software Applications. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3491102.3501895
- Mayer et al. (2018) Mikaël Mayer, Viktor Kuncak, and Ravi Chugh. 2018. Bidirectional Evaluation with Direct Manipulation. Implentation for: Bidirectional Evaluation with Direct Manipulation 2, OOPSLA (Oct. 2018), 127:1–127:28. https://doi.org/10.1145/3276497
- Meyer et al. (2023) Jesse G. Meyer, Ryan J. Urbanowicz, Patrick C. N. Martin, Karen O’Connor, Ruowang Li, Pei-Chen Peng, Tiffani J. Bright, Nicholas Tatonetti, Kyoung Jae Won, Graciela Gonzalez-Hernandez, and Jason H. Moore. 2023. ChatGPT and Large Language Models in Academia: Opportunities and Challenges. BioData Mining 16, 1 (July 2023), 20. https://doi.org/10.1186/s13040-023-00339-9
- Morris et al. (2024) Meredith Ringel Morris, Jascha Sohl-dickstein, Noah Fiedel, Tris Warkentin, Allan Dafoe, Aleksandra Faust, Clement Farabet, and Shane Legg. 2024. Levels of AGI for Operationalizing Progress on the Path to AGI. https://doi.org/10.48550/arXiv.2311.02462 arXiv:2311.02462 [cs]
- Munroe (2009) Randall Munroe. 2009. Movie Narrative Charts. https://xkcd.com/657/.
- Myers (1986) B. A. Myers. 1986. Visual Programming, Programming by Example, and Program Visualization: A Taxonomy. SIGCHI Bull. 17, 4 (April 1986), 59–66. https://doi.org/10.1145/22339.22349
- Neuwirth and Kaufer (1989) C. M. Neuwirth and D. S. Kaufer. 1989. The Role of External Representation in the Writing Process: Implications for the Design of Hypertext-Based Writing Tools. In Proceedings of the Second Annual ACM Conference on Hypertext - HYPERTEXT ’89. ACM Press, Pittsburgh, Pennsylvania, United States, 319–341. https://doi.org/10.1145/74224.74250
- NextUI (2021) nextui-org 2021. NextUI. nextui-org. https://nextui.org/
- OpenAI (2023) OpenAI. 2023. OpenAI Node API Library. OpenAI.
- Paivio (1990) Allan Paivio. 1990. Mental Representations: A Dual Coding Approach. Oxford University Press.
- Pope (1998) Thomas Pope. 1998. Good scripts, bad scripts: learning the craft of screenwriting through 25 of the best and worst films in history. Three Rivers Press.
- Qiang and Bingjie (2016) Lu Qiang and Chai Bingjie. 2016. StoryCake: A Hierarchical Plot Visualization Method for Storytelling in Polar Coordinates. In 2016 International Conference on Cyberworlds (CW). IEEE, Chongqing, China, 211–218. https://doi.org/10.1109/CW.2016.43
- React (2013) Meta 2013. React – A JavaScript Library for Building User Interfaces. Meta. https://reactjs.org/
- Reagan et al. (2016) Andrew J. Reagan, Lewis Mitchell, Dilan Kiley, Christopher M. Danforth, and Peter Sheridan Dodds. 2016. The Emotional Arcs of Stories Are Dominated by Six Basic Shapes. EPJ Data Science 5, 1 (Dec. 2016), 1–12. https://doi.org/10.1140/epjds/s13688-016-0093-1
- Reza et al. (2024) Mohi Reza, Nathan Laundry, Ilya Musabirov, Peter Dushniku, Zhi Yuan ”Michael” Yu, Kashish Mittal, Tovi Grossman, Michael Liut, Anastasia Kuzminykh, and Joseph Jay Williams. 2024. ABScribe: Rapid Exploration & Organization of Multiple Writing Variations in Human-AI Co-Writing Tasks Using Large Language Models. https://doi.org/10.1145/3613904.3641899 arXiv:2310.00117 [cs]
- SaveTheCat (2014) Blake Snyder Enterprises 2014. Save The Cat! Story Structure Software. Blake Snyder Enterprises. https://savethecat.com/save-the-cat-software
- Shneiderman (1983) Shneiderman. 1983. Direct Manipulation: A Step Beyond Programming Languages. Computer 16, 8 (Aug. 1983), 57–69. https://doi.org/10.1109/MC.1983.1654471
- Shneiderman (1982) Ben Shneiderman. 1982. The Future of Interactive Systems and the Emergence of Direct Manipulation†. Behaviour & Information Technology 1, 3 (July 1982), 237–256. https://doi.org/10.1080/01449298208914450
- Shneiderman et al. (2006) Ben Shneiderman, Gerhard Fischer, Mary Czerwinski, Mitch Resnick, Brad Myers, Linda Candy, Ernest Edmonds, Mike Eisenberg, Elisa Giaccardi, Tom Hewett, Pamela Jennings, Bill Kules, Kumiyo Nakakoji, Jay Nunamaker, Randy Pausch, Ted Selker, Elisabeth Sylvan, and Michael Terry. 2006. Creativity Support Tools: Report From a U.S. National Science Foundation Sponsored Workshop. International Journal of Human–Computer Interaction 20, 2 (May 2006), 61–77. https://doi.org/10.1207/s15327590ijhc2002_1
- SlateJS (2016) Slate.js 2016. Slate.js. Slate.js. https://www.slatejs.org
- Suh et al. (2024) Sangho Suh, Meng Chen, Bryan Min, Toby Jia-Jun Li, and Haijun Xia. 2024. Luminate: Structured Generation and Exploration of Design Space with Large Language Models for Human-AI Co-Creation. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 644, 26 pages. https://doi.org/10.1145/3613904.3642400
- Tanahashi and Kwan-Liu Ma (2012) Y. Tanahashi and Kwan-Liu Ma. 2012. Design Considerations for Optimizing Storyline Visualizations. IEEE Transactions on Visualization and Computer Graphics 18, 12 (Dec. 2012), 2679–2688. https://doi.org/10.1109/TVCG.2012.212
- Watson et al. (2019) Katie Watson, Samuel S. Sohn, Sasha Schriber, Markus Gross, Carlos Manuel Muniz, and Mubbasir Kapadia. 2019. StoryPrint: An Interactive Visualization of Stories. In Proceedings of the 24th International Conference on Intelligent User Interfaces (IUI ’19). Association for Computing Machinery, New York, NY, USA, 303–311. https://doi.org/10.1145/3301275.3302302
- Weidinger et al. (2021) Laura Weidinger, John Mellor, Maribeth Rauh, Conor Griffin, Jonathan Uesato, Po-Sen Huang, Myra Cheng, Mia Glaese, Borja Balle, Atoosa Kasirzadeh, Zac Kenton, Sasha Brown, Will Hawkins, Tom Stepleton, Courtney Biles, Abeba Birhane, Julia Haas, Laura Rimell, Lisa Anne Hendricks, William Isaac, Sean Legassick, Geoffrey Irving, and Iason Gabriel. 2021. Ethical and Social Risks of Harm from Language Models. https://doi.org/10.48550/arXiv.2112.04359 arXiv:2112.04359 [cs]
- Williams ([n. d.]) Eric Williams. [n. d.]. How to View and Appreciate Great Movies. https://www.thegreatcourses.com/courses/how-to-view-and-appreciate-great-movies
- Yuan et al. (2022) Ann Yuan, Andy Coenen, Emily Reif, and Daphne Ippolito. 2022. Wordcraft: Story Writing With Large Language Models. In Proceedings of the 27th International Conference on Intelligent User Interfaces (IUI ’22). Association for Computing Machinery, New York, NY, USA, 841–852. https://doi.org/10.1145/3490099.3511105
- Zhang et al. (2023) Zheng Zhang, Jie Gao, Ranjodh Singh Dhaliwal, and Toby Jia-Jun Li. 2023. VISAR: A Human-AI Argumentative Writing Assistant with Visual Programming and Rapid Draft Prototyping. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. ACM, San Francisco CA USA, 1–30. https://doi.org/10.1145/3586183.3606800
- Zhu (2013) Xiaojin Zhu. 2013. Persistent Homology: An Introduction and a New Text Representation for Natural Language Processing. In IJCAI International Joint Conference on Artificial Intelligence. 1953–1959.
- Zhu and Kolassa (2018) Yaqian Zhu and John Kolassa. 2018. Assessing and Comparing the Accuracy of Various Bootstrap Methods. Communications in Statistics - Simulation and Computation 47, 8 (Sept. 2018), 2436–2453. https://doi.org/10.1080/03610918.2017.1348516

### v2

@@
-Visual Writing: Writing by Manipulating Visual Representations of Stories
+Visual Story-Writing: Writing by Manipulating Visual Representations of Stories
 Abstract.
-We introduce “visual writing”, an approach to writing stories by manipulating visuals instead of words.
-Visual writing relies on editable visual representations of time, entities, events, and locations to offer representations more suited to specific editing tasks.
-We propose a taxonomy for these representations and implement a prototype software supporting the visual writing workflow.
-The system allows writers to edit the story by alternating between modifying the text and manipulating visual representations to edit entities, actions, locations, and order of events.
-We evaluate this workflow with eight creative writers and find visual writing can help find specific passages, keep track of story elements, specify edits, and explore story variations in a way that encourages creativity.
+We define “visual story-writing” as using visual representations of story elements to support writing and revising narrative texts.
+To demonstrate this approach, we developed a text editor that automatically visualizes a graph of entity interactions, movement between locations, and a timeline of story events.
+Interacting with these visualizations results in suggested text edits: for example, connecting two characters in the graph creates an interaction between them, moving an entity updates their described location, and rearranging events on the timeline reorganizes the narrative sequence.
+Through two user studies on narrative text editing and writing, we found that visuals supported participants in planning high-level revisions, tracking story elements, and exploring story variations in ways that encourage creativity.
+Broadly, our work lays the foundation for writing support, not just through words, but also visuals.
 1. Introduction
-Writing stories and exploring alternative narratives can be challenging.
-Although using large language models (LLMs) can help writers (Chakrabarty et al., 2024), it remains difficult for users to formulate their intent using words (Dang et al., 2022b) and review the generated text to ensure it aligns with their intent (Masson et al., 2024).
+Creative writers manage many moving parts, from character arcs and causal chains to spatial coherence and narrative timing (Bal, 1997; Norton, 2011).
+Keeping track of everything is challenging, and it gets even more complicated when experimenting with story ideas (Zhao et al., 2025).
+Modifications require careful planning and multiple rounds of edits.
 Consider a task as simple as changing the location of a character in a story: the cat no longer goes to the barn but instead wanders about the lake.
-Using “replace all” instances of ‘barn’ is insufficient because substituting a location comes with implications.
-For example, descriptions and actions pertaining to the barn would need to be updated to reflect the new context of being by the lake.
-Further, not all instances necessarily need to be impacted.
-The fact that the cat no longer visits the barn should not result in the barn being completely removed from the story.
-Instead, writers must carefully plan to find all and only the references to the location associated with the character’s change of action and make tedious revisions to the text to maintain spatial and semantic consistencies.
-This busywork is prone to errors, and writers often resort to creating external documents such as maps, spreadsheets, and timelines that they painstakingly update to keep track of the characters, events, and locations (Marshall, 2014; Ackerman, 2017; Neuwirth and Kaufer, 1989; SaveTheCat, 2014).
-While LLMs could help with this busywork, expressing distinct changes can be challenging with prompts (Masson et al., 2024), and the extent and suitability of the changes made by the generative model could be difficult to verify.
-The challenge stems from the mismatch between the reasoning and representation spaces: in our above example, the writer performs spatial reasoning using a textual representation.
-A spatial representation, such as a map of the story world, would be more suited for such a task, where changing the location of the character becomes as simple as moving its position on the map.
-In text, edits relevant to space and locations in the story are comparatively challenging to accomplish, which can be a deterrent for writers wanting to change the narrative and explore variations.
-And this same argument applies to all other aspects of a story such as time, events, and characters, for which we posit superior representations can be designed.
-We define visual writing as the use of visual representations of story elements as an input medium for expressing writing intents.
-For each story element, we explore the use of a visual representation that conceptually aligns with that element.
-Depending on their editing task, users can decide to edit the text of the story, in which case the visual representations are updated to match the new text.
-Alternatively, they can manipulate the visual representations through direct manipulation to suggest changes to the text and rapidly explore variations.
-We discuss a taxonomy of these visual representations, drawing on the story elements framework from narratology (§3).
-In a prototype software (§4), we implement three of these visual representations: (1) a diagrammatic view of the entities in the story and their relationships, allowing to add, remove, and edit characters and add new actions between them (Figure 1.c); (2) a spatial view of the locations in the story and the locations of the entities, allowing to add new locations and move entities (Figure 1.d); and (3) a timeline of the events in the story, allowing to reorder the events, quickly find specific scenes, and precisely modify the entities and locations of the selected scenes by leveraging the other two views (Figure 1.b).
-We validate this approach in a user study with eight creative writers (§5).
-Results show participants appreciated the visual representations to locate scenes and keep track of story elements while writing.
-They were successful in expressing their editing intents through direct manipulations of the visual elements.
-Furthermore, participants appreciated this approach to edit text from an exploration and creativity perspective.
-This points to the potential of visual representations for writing and calls for more work in that direction.
+Simply replacing all instances of “barn” would cause inconsistencies unless descriptions and actions pertaining to the barn are also updated to reflect the new context of the lake.
+Moreover, the fact that the cat wanders about the lake does not mean that the barn needs to be removed entirely from the story.
+Instead, before revising the text, writers must identify references to the barn and tediously review the other characters’ locations to maintain spatial and semantic consistency.
+This busywork is prone to errors, and writers often resort to creating external documents, such as maps, spreadsheets, and timelines, that they painstakingly update to keep track of characters, events, and locations (Zhao et al., 2025; Ackerman, 2017; Neuwirth and Kaufer, 1989; SaveTheCat, 2014; Articy, 2014).
+Large language models (LLMs) could help, but expressing specific intents with a prompt is difficult (Masson et al., 2024; Zamfirescu-Pereira et al., 2023), and trying to explain which barn and which cat to edit will leave much room for misinterpretation.
+Further, the suitability of the changes made by the generative model could be difficult to verify.
+This challenge stems from the mismatch between the reasoning and representation spaces: in our barn example, the writer does spatial reasoning with a textual representation.
+A better approach would be to use a spatial representation, such as a map of the story world, where changing the location of a character becomes as simple as moving it on the map.
+This same argument applies to all other aspects of a story, such as time, events, and characters, for which we posit superior representations can be designed.
+Therefore, we propose using representations that match the reasoning space.
+We define visual story-writing as the use of visual representations of story elements as both a reviewing tool and an input medium for expressing writing intents (Figure 1).
+This complements the writing workflow of writers with various experience levels.
+For example, an experienced writer may prefer editing the story’s text, in which case the visual representations updates, offering a visualization to help the writer track story elements.
+Alternatively, a beginner writer could manipulate the visual representations through direct manipulation to explore story variations.
+We articulate a framework grounded in narratology to design new visual representations of stories (§3).
+Applying our framework, we explore the potential of visual story-writing through a prototype software (§4) with three illustrative visual representations: (1) a diagrammatic view of the entities in the story and their relationships, allowing to add, remove, and edit characters and add new actions between them; (2) a spatial view of the entities in the story, allowing to add new locations and move entities; and (3) a timeline of the events in the story, allowing to reorder the narrative, quickly find specific scenes, and precisely modify the entities and locations of the selected scenes by leveraging the other views.
+We validate this approach through two user studies looking at the different components of the writing process (Flower and Hayes, 1981).
+Results from our first study revealed that visuals helped plan high-level revisions, search, and promote critical reflection over having only text.
+Results from our second study show the potential of visual story-writing to express editing intents and help writers explore story variations in a way that supports creative expression.
+This points to the potential of visual story-writing and encourages further exploration.
 2. Related Work
-Visual writing draws inspiration from work on visualizing stories and efforts to give alternative representations to edit content. Below, we review the literature on these topics.
+Visual story-writing draws inspiration from work on visualizing stories and efforts to give alternative representations to edit content.
 2.1. Visualizing Stories
-Visual representations of stories and story patterns have a long-standing presence in writing, with staple examples including Freytag’s pyramid story (Freytag, 1895) and abstract diagrams capturing the “shape of the story” (Pope, 1998; Comberg, [n.
+Visual representations of stories and story patterns have a long-standing presence in writing, with staple examples including Freytag’s pyramid story (Freytag, 2004) and abstract diagrams capturing the “shape of the story” (Pope, 1998; Comberg, [n.
 d.]; Reagan et al., 2016).
-Such visual representations are not only useful for teaching the craft of storytelling, they are also considered as useful aids to support the writing process.
+Such visual representations are not only useful for teaching the craft of storytelling but are also considered useful aids to support the writing process.
 Referring to the visual representations, screenplay writer E.
-Williams says: “I find story shape to be a great tool […] to make sure my story holds together” (Williams, [n.
-d.]).
-There exist many techniques to visualize stories, spanning from classical charts like word clouds, to more artistic visualizations like calligrams (Brath, 2021).
-Often, these visualizations summarize entire stories in an attempt to highlight a specific facet (e.g., emotions (Maharjan et al., 2018) or topic recurrence (Zhu, 2013)).
+Williams says: “I find story shape to be a great tool […] to make sure my story holds together” (Williams, 2018).
+There exist many techniques to visualize stories, spanning from classical charts, like word clouds, to more artistic visualizations like calligrams (Brath, 2021).
+Often, these visualizations attempt to highlight a specific facet of the story (e.g., emotions (Maharjan et al., 2018), motion (Chung et al., 2025), fortune (Comberg, [n.
+d.]; Chung et al., 2022), or topic recurrence (Zhu, 2013)).
 For example, Storyline visualizations (Munroe, 2009; Tanahashi and Kwan-Liu Ma, 2012) depict the evolution of the relationships between characters through timelines.
 StoryCake (Qiang and Bingjie, 2016) uses a hierarchical plot to visualize the structure of nonlinear stories.
-Story Curves (Kim et al., 2018) visualizes non-linear narratives by showing both the order in which the events are told and the order of the underlying story.
+Story Curves (Kim et al., 2018) visualizes non-linear narratives by showing both the order in which the events are told and the story’s chronological order.
 StoryPrint (Watson et al., 2019) shows the scenes and character emotions through circular timelines.
 Each of these visual representations presents the ability to support readers, writers, and analysts alike in appreciating or even analyzing the stories from a specific lens.
-Other story visualizations have been used in tools to help during script and story writing.
-For example, CARDINAL (Marti et al., 2018) is a tool to help screenwriters visualize their script through different views, including a visualization of the interactions between characters and an animated 2D and 3D views to play the scene and visualize locations and movements of the characters.
+Other story visualizations have been used in tools to help with script and story writing.
+For example, CARDINAL (Marti et al., 2018) helps screenwriters visualize their script through different views, including a visualization of the interactions between characters and animated 2D and 3D views to playout the scene and replay character movements.
 Similarly, Portrayal (Hoque et al., 2023) attempts to help writers develop their characters by extracting and presenting information related to the characters in interactive visualizations.
-Our work explores the potential of expanding the role of such visualizations to also be manipulable.
-That is, by mapping visualization elements directly to entities, events, and locations in the text story, we explore using visualizations as the medium to modify the story.
-This approach is motivated by the potential of visual representations to express intent, which we discuss next.
-2.2. Editing Content Through Alternative Representations
-Programming might be the domain that most embraced the use of alternative representations.
-Approaches include visual programming (Myers, 1986; Burnett and McIntyre, 1995), which allows generating programs by manipulating elements in more than one dimension, code projections (Gobert and Beaudouin-Lafon, 2023; Ko and Myers, 2006) which augment programming languages with tools and widgets that are alternative representations of fragments of code, and bi-directional programming (Mayer et al., 2018; Hempel et al., 2019) where the output of a program can be manipulated to edit the program itself.
-Similar to our work, the motivation behind these approaches is that some tasks are easier done using different representations: it is easier to find a colour using a colour picker than by typing its hexadecimal code, and it is easier to move an object on the screen through direct manipulation (Shneiderman, 1983) than by typing its coordinates.
-In the creative writing domain, this idea of manipulating projections or visual representations to edit stories is largely untapped, which is surprising given the reliance on visual representations to capture story concepts (Freytag, 1895; Comberg, [n.
-d.]).
-Instead, most writing support tools ask users to manipulate text using text, typically through prompts or suggestions (Calderwood et al., 2020; Buschek et al., 2021; Yuan et al., 2022; Dang et al., 2023; Lee et al., 2024; Reza et al., 2024).
-There are some exceptions, such as the proposal of Kempen et al.
-(1987) to represent and manipulate text in different ways, such as manipulating the syntax tree to allow restructuring sentences.
-Similarly, Arnold et al.
-(2021) suggested that words could be directly manipulated to be reordered.
+Our work seeks to inform the design of story visualizations by proposing a new framework.
+This framework helps generate representations that might serve as a medium to edit the story, a topic we discuss next.
+2.2. Editing Using Alternative Representations
+Visualizations are increasingly used to not only encode data but also collect and modify new data (Bressa et al., 2024).
+In our work, visual representations can be edited to modify the narrative text in addition to exploring and analyzing it, making it an input visualization.
+One domain which has embraced the use of input visualizations is computer programming.
+For example, visual programming (Myers, 1986; Burnett and McIntyre, 1995) allows generating programs by manipulating elements in more than one dimension, code projections (Gobert and Beaudouin-Lafon, 2023; Ko and Myers, 2006) augments programming languages with tools and widgets that are alternative representations of fragments of code, and bi-directional programming (Mayer et al., 2018; Hempel et al., 2019) allows manipulating the output of a program to edit the program itself.
+Similar to our work, the motivation behind these approaches is that some tasks are easier done using different representations: it is easier to find a colour using a colour picker than by typing its hexadecimal code, and it is easier to move an object through direct manipulation (Shneiderman, 1983) than by typing coordinates.
+In the creative writing domain, the idea of manipulating projections or visual representations to edit stories is largely untapped despite the reliance on visual representations to capture story concepts (Freytag, 2004; Comberg, [n.
+d.]; Zhao et al., 2025).
+Instead, most writing support tools ask users to manipulate narratives using text, typically through prompts or suggestions (Calderwood et al., 2020; Buschek et al., 2021; Yuan et al., 2022; Dang et al., 2023; Lee et al., 2024; Reza et al., 2024).
+Some exceptions include the proposal of Kempen et al.
+(1987) to manipulate the syntax tree of a sentence to restructure it or the proposal to directly drag the words to reorder them (Arnold et al., 2021).
 Beyond syntactic level modifications, Dang et al.
-(2022a) proposed a system that allows reorganizing, merging, and removing paragraphs by manipulating textual summaries instead of the text.
-In the same vein, VISAR (Zhang et al., 2023) automatically generates an interactive visual outline of an argumentative text, which can be edited to change the structure of the argument.
-Finally, TaleBrush (Chung et al., 2022) leverages a canvas where sketching allows modulating the fortune of characters in a story.
-Our work contributes to this line of research by formalizing the concept of visual writing.
+(2022) proposed a system that allows reorganizing, merging, and removing paragraphs by manipulating textual summaries.
+Moving even further from text as an input medium, the machine learning task of “visual storytelling” consists in generating a story from a sequence of images (Huang et al., 2016; Hsu et al., 2019b; Wang et al., 2018; Hsu et al., 2019a; Bensaid et al., 2021).
+Similarly, TaleBrush (Chung et al., 2022) leverages a canvas where sketching allows modulating the fortune of characters in a story, whereas Textoshop (Masson et al., 2025) offers features inspired by drawing software such as tone pickers and text layers.
+Closest to our approach are systems that provide an editable alternative view of the text.
+For example, VISAR (Zhang et al., 2023) generates a visual outline of an argumentative text, which can be edited to change the structure of the argument.
+Toyteller (Chung et al., 2025) generates story text based on character motions, as if the characters were toys.
+Similarly, XCreation (Yan et al., 2023) supports creativity and cross-modality edits through the use of clip-arts and a relationship diagram as an intermediate representation between text and images.
+Our work contributes to this line of research by formalizing visual story-writing.
 The concept is rooted in the same principles as visual programming and code projections, positing that elements of a story become easier to understand and manipulate when presented in alternative forms.
-Specifically, we design visualizations grounded in narratology to explore this approach.
-3. Taxonomy of Visual Representations
-Narratology separates the chronological sequence of the story (fabula) from the narrative sequence (syuzhet) of the story.
-At its core, a fabula consists of four main story elements: events, actors111In narratology, actors refer to “agents that perform actions.
+Specifically, we propose a framework grounded in structural narratology theories to inform the design of visualizations and extend the concept to all story elements, including events, like XCreation (Yan et al., 2023), but also space and time.
+3. A Framework for Story Constructs
+To inform the design of visual representations that support the visual story-writing workflow, we articulate a generative framework.
+Our framework builds upon narrative theories by structuralists like Gérard Genette (Genette et al., 1990) and Mieke Bal (Bal, 1997, 2021).
+Elements of stories outlined by structuralists (Bal, 1997, 2021) match what we observed in existing story visualizations, such as organizing story elements based on time (Munroe, 2009; Watson et al., 2019), space (Marti et al., 2018), or a combination thereof (Hulstein et al., 2023).
+Story elements do not work in isolation.
+Rather, many important aspects of stories stem from the intricate weaving of these elements, such as the movement of characters across locations over time.
+We propose an operator mechanism to combine elements into such meaningful composites.
+Our operators are applied sequentially to story elements to form a story construct that can then be represented and manipulated.
+This framework can describe story constructs in existing visualizations and also generate new ones.
+position to place elements spatially;
+associate to add new elements;
+connect to add edges; and
+unfold to duplicate and organize elements.3.1. Story Elements
+Story elements are the fundamental building blocks of story constructs found in story visualizations.
+Within narratology, the early Russian formalists separated the chronological order of the story (fabula) from the order of the plot or how it is represented to readers (syuzhet) (Liveley, 2021; Bal, 1997).
+Structuralists like Genette and Bal borrowed these concepts to analyze narrative structurally.
+Bal (Bal, 1997, 2021) describes a fabula, the factual events within a story, to consist of four main elements: actors111Actors refer to “agents that perform actions.
 They are not necessarily human” (Bal, 1997).
-To avoid confusion, we use the term entity instead throughout this paper., time, and location (Bal, 1997).
-Similarly, we define visual writing as the use of visuals that represent and manipulate these four elements.
-In their simplest form, visual representations can depict elements one at a time.
-For example, a list of locations would allow adding, removing, and editing the characteristics of each location.
-And a list of entities, similar to the character sheets writers often create to plan their stories (SaveTheCat, 2014), would allow adding and removing entities.
-However, many important aspects of stories are encoded through the interaction between elements (e.g., the time of an event or the location of an entity).
-Below, using the narratology lens, we discuss a non-exhaustive taxonomy of possible visual representations as an initial guide to inform the design of tools supporting visual writing.
-Spatial representations depict elements based on their location.
-For example, a spatial representation could be a map with dots, where each dot represents where events happened and where entities are located.
-Or it could be more abstract and only represent distances between elements.
-Spatial representations are used by systems like CARDINAL (Marti et al., 2018) to show the relative positioning of characters within a scene.
-Event-driven representations depict elements based on the events they relate to.
-For example, a node graph could be an event-driven representation if each node represents an event and the edges encode another element such as the entities (in which case, connected nodes would represent events affecting the same entity) or time (in which case the directed edges would indicate the flow of the events).
-Some examples of event-driven representations are lists of outline points and event graphs.
-Entity-driven representations depict elements based on the entities they relate to.
-For example, a node graph could be an entity-driven representation if each node represents an entity and the edges encode another element such as the events (in which case, each edge is an event connecting the entities involved in the event) or time (in which case, the directed edges indicate the order in which the entities intervene in the story).
-Entity-driven representations, such as character sheets and family trees, are often created by writers to keep track of character details (SaveTheCat, 2014).
-Temporal representations depict elements based on the time they happen in the fictional world of the story or in the order they are presented in the text.
-For example, a timeline could arrange elements such as events (in which case, the events would be positioned based on when they occur) or entities (in which case, entities would be positioned based on when they appear or when they perform an action).
-Temporal representations have been more commonly used by previous work on visualizing stories such as Storylines (Munroe, 2009), StoryCake (Qiang and Bingjie, 2016), and StoryPrint (Watson et al., 2019).
-The design space for the above categories is vast, and there exist many visualizations to bear inspiration from.
-Our goal is not to argue which specific design is best—we leave it to future work to explore different variations within these categories (and beyond).
-Instead, our work focuses on the potential of visual writing, i.e.
-the use of alternative representation of stories, as an input medium for expressing intent.
-4. The Visual Writing Prototype
-To test the potential of visual writing, we developed a prototype system, shown in Figure 1, implementing three visual representations: a diagrammatic view of the entities and actions (Entity–Event), a spatial view of the locations and entities (Location–Entity), and a timeline of the events (Time–Event).
-The focus of the prototype was to help writers explore variations and perform spatial and temporal edits via visual writing.
-As such, the system allowed specifying edits through direct manipulations of the graphical elements (Shneiderman, 1982; Hutchins et al., 1985) and followed the principles to design creativity support tools (Shneiderman et al., 2006).
-Below, we detail the design and features of the system.
-All the examples and screenshots were taken using the system on the first three paragraphs of the story Alice’s Adventures in Wonderland by Lewis Carroll.
+To avoid confusion, we use the term entity instead in our prototype system., time, location, and events.
+The syuzhet, or how the story is told, also consists of four elements: characters, temporality, space, and focalization.
+See table 1 for examples.
+| Element of Fabula | Element of Syuzhet |
+| Actor (e.g., hero, villain) | Character (e.g., Alice) |
+| Location (e.g., Alice’s house) | Space (e.g., home, eerie) |
+| Time (chronological) | Temporality (narrated order) |
+| Event (what happens) | Focalization (point of view) |
+Although elements of fabula and syuzhet may seem similar, they refer to distinct concepts.
+For instance, characters are concrete entities (e.g., Alice, the white rabbit), whereas actors are abstract roles or functions characters can take (e.g., hero, villain).
+Time corresponds to the chronological timeline of events, whereas temporality corresponds to how those events are narrated (e.g., through flashbacks, ellipsis).
+And locations are concrete, physical places (e.g., Alice’s house, a forest), whereas spaces refer to the setting and the narrated representation of these locations (e.g., home, eerie).
+We describe visual story-writing as the use of visuals that represent and manipulate story constructs, i.e.
+composites of these eight elements.
+3.2. Operators
+Our operators are functions that combine story elements (operands) into higher-level story constructs.
+They can be chained to create more elaborate constructs.
+Based on existing story visualizations, we derived four binary operators acting on operands and 222Note that our operators are noncommutative.
+Thus, while is always one of the eight story elements, can be a composite of elements resulting from a chain of operations.:
+position (
+) places elements of based on the story element (Figure 2.
+). For example, positioning time by locations results in a map of the movements in the story. Note that the only valid operands are location and space. -
+associate (
+) adds the elements of and associates them to the elements of (Figure 2.
+For example, associating time with focalization creates a timeline of point of view. -
+connect (
+) adds edges between elements of , according to that of (Figure 2.
+). For example, connecting characters by events creates a graph of character interactions. -
+unfold (
+) duplicates and organizes elements of based on that of (Figure 2.
+). This essentially duplicates the elements according to each element of . For example, unfolding locations by characters creates a list of locations visited by each character.
+3.3. Describing Existing Story Visualizations
+We test our framework by describing existing story visualizations.
+Storyline (Munroe, 2009) shows character interactions: start with time, unfold by characters, and connect by events.
+StoryCurve (Kim et al., 2018) shows nonlinear narratives: start with time, unfold by temporality to visualize the non-linear narrative, associate with locations (using colours) and associate with events.
+StoryPrint (Watson et al., 2019) shows scenes, character presence and emotions: start with time, unfold by characters, and associate with events or other elements (StoryPrint also encodes emotions, which goes beyond the scope of structural narratology).
+Geo-Storylines (Hulstein et al., 2023) shows locations of characters over time: start with Storyline, position by locations.
+The map glyph variation: start with locations, position by location, unfold by time, and associate with characters.
+“All fights from Dragon Ball Z”333A non-academic example (Cinnamon, 2017) shows the battles between characters within a saga: start with events (only the fights), unfold by time, and connect by characters.
+3.4. Generating New Story Constructs
+Figure 2 shows examples of chaining operators to create more complex story constructs.
+The number of unique combinations of operators and elements is vast.
+Our goal is not to argue which specific story construct is optimal for writers nor which visual design is best to represent these constructs—we leave it to future work to explore variations within these categories (and beyond).
+Instead, our work focuses on the potential of visual story-writing, i.e.
+the use of alternative representations of stories, as an input medium for expressing intent.
+As such, we designed three simple visual representations of story constructs interconnected through interactivity (North and Shneiderman, 2001; Wang Baldonado et al., 2000): a timeline of entities and events (start with time, unfold by characters and connect by events); a graph of actions and entities (start with characters and connect by events).
+And a view of the locations (start with characters and position by locations).
+There are other important components to a story, such as character traits and emotions, but in Bal’s narratology model (Bal, 1997, 2021), these are properties of the story elements.
+One way to represent them within the framework is as an attribute of a story element.
+3.5. Recommending Interactions and Visuals
+The framework can be further expanded to prescribe interactions and visual representations.
+For example, applying the position operator suggests that the placed elements can be dragged to express a change of location; the associate operator suggests a detail-on-demand type of interaction to view the associated elements; the connect operator suggests an interaction to remove and create links between elements; and the unfold operator suggests elements can be dragged to be reordered or reassigned.
+Similarly, the framework could also prescribe visual representations by mapping operations to specific visual channels.
+4. The Visual Story-Writing Prototype
+To test the potential of visual story-writing, we developed a prototype system, shown in Figure 3, with three visual representations informed by our framework: a diagrammatic view of the entities and actions, a spatial view of the locations and entities, and a timeline of the events (section 3.4).
+The focus of the prototype was to help writers explore variations and perform spatial or temporal edits via visual story-writing.
+As such, the system allowed specifying edits through direct manipulations of the graphical elements (Shneiderman, 1982; Hutchins et al., 1985) and followed the design principles of creativity support tools to support exploration (Shneiderman et al., 2006).
+Below, we detail the system design and features (marked with
+when the feature helps review the story,
+when it helps edit the story, and
+when it helps for both editing and reviewing). All screenshots were taken using the system on the first three paragraphs of the story Alice’s Adventures in Wonderland by Lewis Carroll.
 4.1. Entities and Actions View
-This event-driven representation shows all the entities in the story as nodes with their name and an emoji representing them (Figure 1.c).
+This representation shows all the entities in the story as nodes with their name and an emoji representing them (Figure 3.b).
 The actions between the entities are represented as directed edges with a label describing the action in one or two words.
 By default, the nodes are arranged using a force layout to prevent overlaps, but users can rearrange the position of the entities by dragging them.
-4.1.1. Editing the Traits of an Entity
-Selecting an entity expands its node to reveal traits related to the entity similar to a list of personality traits which would be present on a character sheet (e.g., “curious”, “adventurous”, “chatty”).
-The intensity of each trait is rated from 1 to 10.
-This intensity can be changed by moving the slider beside the trait.
-This results in editing the text of the story to reflect the new intensity of the trait (Figure 2.a).
-4.1.2. Adding and Removing an Entity
-Adding an entity is done by double-clicking anywhere on the canvas which opens a text input box to indicate the name of the new entity.
-Once the name is entered, the entity is created at the location of the click.
-Removing an entity is done by selecting the entity and then pressing “delete” or “backspace” on the keyboard.
+4.1.1.
+Editing the Traits of an Entity
+Selecting an entity opens a menu with traits similar to a list of personality traits which would be present on a character sheet (e.g., “curious”, “adventurous”, “chatty”).
+The intensity of each trait is rated on a scale of 1 to 10.
+Moving the slider beneath the trait results in editing the text of the story to reflect the new intensity of the trait (Figure 4.a).
+4.1.2.
+Adding and Removing an Entity
+Double-clicking on the canvas opens a text input box, which, once filled, creates an entity with the given name at the location of the mouse pointer.
+Selecting an entity and then pressing “delete” or “backspace” on the keyboard removes the entity.
 This results in updating the text of the story so that the entity is effectively removed (Figure 4.b).
-4.1.3. Adding and Removing Actions between Entities
-Adding an action is done by creating an edge between two nodes.
+4.1.3.
+Adding and Removing Actions between Entities
+Creating an edge between two nodes opens a text input field to enter a new action.
 The direction of the edge indicates the source and target of the action.
-An action can also be created for the same character by connecting a node to itself.
+Connecting a node to itself means the action is for the same character.
 This helps represent actions done by an entity with no target (e.g., walking or thinking).
-An action can be edited at all times by double-clicking its label and changing its description.
-Similarly, an action can be removed by being selected and then pressing “delete” or “backspace”.
+Double-clicking the label of an action edits it.
+Similarly, selecting an action and pressing “delete” or “backspace” removes it.
 As soon as the edge is created, edited, or deleted, the story text is updated to reflect the change (Figure 4.c).
+4.1.4.
+Overlapping Edges and Animations
+To declutter the view, overlapping edges are grouped so that only the first action is shown alongside a counter.
+Users can cycle through the events by clicking the next and previous buttons.
+Alternatively, hovering over an entity shows animated dots coming and going from the entity, with the name of the action beside them.
 4.2. Locations and Entities View
-The location view is a spatial representation accessed through a tab (Figure 1.d).
-It shows all the locations in the story as nodes with their name and an emoji representing them.
-Similar to the entities and actions view (Section 4.1), the view shows all entities as smaller nodes, but their position is conditioned by their location in the story.
-Entities that go to different locations in the story are duplicated to be represented in all of their locations.
-4.2.1. Moving an Entity
+The location view is a spatial representation accessed through a tab (Figure 3.c).
+It displays all the locations in the story as nodes with their name and an emoji representing them.
+Similar to the entities and actions view (Section 4.1), the view shows all entities as smaller nodes, but their position depends on their order of appearance in the story.
+Entities that go to different locations in the story are duplicated to be represented in all their locations.
+4.2.1.
+Adding a Location
+Double-clicking anywhere on the canvas opens a text input field to create a new location (Figure 5.a). Once the location is created, entities can be moved to it.
+4.2.2.
+Moving an Entity
 Entities can be dragged and moved around the view.
@@
 Once the entity is moved, the story text is updated to reflect the change in location.
-4.2.2. Adding a Location
-A new location can be added by double-clicking anywhere on the canvas (Figure 3.a).
-This opens a text input field to specify the name of the location.
-Once the location is created, entities can be moved to it.
 4.3. Timeline of Events View
-The timeline view is a temporal representation showing the events in the story represented as vertical lines with the emojis of the entities involved in the event on either side of the line (Figure 1.b).
-These lines are organized one after the other based on the order in which they are presented in the text.
+The timeline view is a temporal representation showing the events in the story represented as vertical lines with the emojis of the entities involved in the event on either side of the line (Figure 3.a).
+These lines are organized one after the other based on the order in which they are presented in the text (temporality).
 The timeline view is always visible.
 It allows selecting one or multiple events, in which case the other views are updated to fade out the entities and locations not involved in the selected events.
-Selecting events also forces future modifications to impact only the selected events.
-4.3.1. Reordering Events
-Events can be dragged to be reorganized temporally.
-When an event is dragged, it slides horizontally and can be placed between other events (Figure 4.b-c).
-Once the event is moved, the story text is updated so that the event happens when indicated in the timeline.
-4.3.2. Selecting Events
+Selecting events also forces subsequent modifications to impact only the selected events.
+4.3.1.
+Selecting Events
 The timeline also allows finding specific scenes and selecting them to allow precise editing.
@@
 Clicking selects the event, and dragging selects multiple events.
-Once one or multiple events are selected, any visual edits made in the views will target these selected events.
-For example, selecting an event and then moving an entity in the location view will update the story text so that the entity moved exactly during the selected event.
+Once events are selected, future visual edits will target these selected events.
+For example, selecting an event and then moving an entity in the location view updates the story text so that the entity moved exactly during the selected event.
+4.3.2.
+Reordering Events
+Dragging an event slides it horizontally (Figure 6.b-c). Once moved, the story text updates so that the event happens when indicated in the timeline.
 4.4. Bi-directional Editor
-The rest of the interface is composed of a text editor on the left, a history tree at the bottom to navigate different versions of the story, and interface buttons at the centre to refresh the visuals or rewrite the story (Figure 1.a).
-4.4.1. Highlighting Visual Elements and Sentences on Hover
-Similar to how hovering over the events in the timeline highlights the sentences, hovering over the sentences in the text editor highlights the corresponding events, entities, and locations in the different views.
-This allows for quickly finding different scenes.
+The rest of the interface consists of a text editor on the left, a history tree at the bottom, and interface buttons at the centre to refresh the visuals or rewrite the story (Figure 3).
+4.4.1.
+Highlighting Visual Elements and Sentences on Hover
+Similar to how hovering over events in the timeline highlights the corresponding sentences, hovering over the sentences in the text editor highlights the corresponding events, entities, and locations in the different views.
+This helps find scenes.
 Similarly, placing the caret in one sentence will act as if the events in that sentence were selected in the timeline.
-As such, future edits to the visual representations will be restricted to that sentence.
-4.4.2. Updating the Visual Representations from the Text
+As such, subsequent edits to the visual representations will be restricted to that sentence.
+4.4.2.
+Updating the Visual Representations from the Text
 When the text is updated manually, the visual representations might become desynchronized, showing the story as it was before the modification.
-In this case, the button to refresh the visuals becomes blue to indicate that the visual model may need to be updated.
-Clicking it results in re-extracting the information from the text and refreshing the views.
-4.4.3. Writing the Story from the Visual Representations
-By clicking the “refresh from visuals” button, the story is rewritten from scratch using only the visual representations as reference.
-As such, the story preserves the same events, locations, and entities, but everything else is rewritten.
-This allows specifying the skeleton of a story and quickly exploring different ways of phrasing the story.
-4.4.4. Track changes
-By default, all changes made by editing the visual representations are tracked, such that removed text is struck in red while additions are shown in blue.
-This allows writers to easily locate and review the changes before accepting (or rejecting) them.
-4.4.5. History Tree
-To make edits easily reversible and facilitate the exploration of alternatives, the system also implements a history tree capable of storing different versions of the story and returning to them at any time.
-Every time a modification happens, be it by editing the visual representation or editing the text of the story, a snapshot of the story and visual representation are added to the tree.
-If the writer selects a previous snapshot, then the story is changed to that snapshot.
-If, after selecting a previous snapshot, a modification is made to the story, then this creates a new branch to the story.
-Then, writers can navigate between the branches and different versions by clicking on the nodes of the history tree.
+In this case, the refresh button becomes blue to indicate that the visual model might be outdated.
+Clicking it re-extracts the information from the text and refreshes the views.
+4.4.3.
+Writing the Story from the Visual Representations
+By clicking the “refresh from visuals” button (arrows between the two views in Figure 3), the story is regenerated from scratch using only the visual representations as reference.
+The regenerated story preserves the same events, locations, and entities, but everything else is rewritten.
+This allows for specifying the skeleton of a story and quickly exploring different ways of phrasing it.
+4.4.4.
+Track changes
+Changes made by editing the visual representations are tracked, such that removed text is struck while additions are highlighted in green.
+This allows writers to locate and review the changes before accepting (or rejecting) them.
+4.4.5.
+History Tree
+To make edits easily reversible and facilitate the exploration of alternatives, the system implements a history tree to store different versions of the story and return to them at any time.
+When a modification happens, be it by editing the visual representation or editing the text, a snapshot of the story and visual representations is added to the tree.
+When the writer selects a previous snapshot, the story is reverted to that snapshot.
+If, after selecting a previous snapshot, a modification is made, then a new branch is created.
+This allows writers to navigate between branches and versions by clicking on the nodes of the history tree.
 4.5. Implementation
-We implemented the visual writing system using TypeScript and React (React, 2013) for the interface, NextUI (NextUI, 2021) for the graphical components, Slate.js (SlateJS, 2016) for the text editor, and the OpenAI library for prompting large language models (OpenAI, 2023).
-Under the hood, modifications to the visual representations are turned into engineered prompts sent to OpenAI’s “GPT4-o” model.
+We implemented the visual story-writing system using TypeScript and React (React, 2013) for the interface, NextUI (NextUI, 2021) for the graphical components, Slate.js (SlateJS, 2016) for the text editor, and the OpenAI library for prompting large language models (OpenAI, 2023).
+Modifications to the visual representations are turned into engineered prompts sent to OpenAI’s “GPT4-o” model.
+A live demo and the source code are accessible online: https://github.com/m-damien/VisualStoryWriting
 4.5.1. Extracting Entities, Locations, and Events
-The visual writing approach requires a robust, exhaustive, and fast extraction of the information needed for the visual representations.
-In our tests, a naive approach consisting of giving a large language model the whole story and asking to extract the necessary information failed at all three of these requirements.
-Instead, we force the model to output JSON with the proper schema to make the extraction robust.
-We check that the schema is correct and prompt the model again if necessary.
-To make the extraction exhaustive, we proceed in multiple steps by first asking the model to extract the entities and locations in the story.
-We then ask the model to extract events between these entities and at those locations.
-Finally, and most importantly, we split the whole story into sentences to make the extraction fast.
-We then prompt the model with the story, but we ask it to extract only the information within the sentence.
-This allows us to execute the prompt requests in parallel.
-As such, a text with 30 sentences will send 30 different smaller prompts to the model.
-This also gives us an indication of the position of the extracted information, as we know that it comes from the sentence being extracted.
+In our tests, the naïve approach of asking an LLM to extract the necessary information results in slow and incomplete extractions.
+Instead, we first ask the model to extract the entities and locations in the story.
+We then split the text into sentences and parallelize requests by giving the model the whole story but asking it to extract only the information within each sentence (section A.1).
+As such, a text with 30 sentences sends 30 smaller prompts in parallel.
+This leads to faster and more exhaustive extractions.
+It also helps associate sentences with events.
+All requests use a structured JSON output.
 4.5.2. Engineered Prompts to Edit the Story
-The system relies on many engineered prompts such as splitting requests to make them faster, giving the model only a subset of the text to force localized edits, and giving the model the previous abstract representation of the story and the new one to help it understand the edits.
-We encourage interested readers to review the source code for more details.
-5. User Study
-We evaluated the visual writing workflow by conducting a user study using our prototype system.
-Specifically, the study tried to answer the following questions: Are the visual representations understood and helpful?
-Can people express their editing intents using visual writing?
-What is the impact of visual writing on exploration and creativity?
-Our study is separated into two parts.
-The first part isolates the use of the three different views in a controlled setting.
-Because our approach is compatible with other ways to edit text (e.g., no assistance or LLMs) and our main interest is to measure the usefulness of visual writing, we decided against comparing our approach to a baseline.
-Instead, we collected qualitative insights through semi-structured interviews and explicitly asked participants to compare visual writing with their own workflow.
-Representations were also evaluated quantitatively using simple usability questionnaires.
-The second part of our study was a free-form creative writing task to let participants appropriate the tool and its features.
-5.1. Participants
-We recruited 8 creative writers from our social network and word of mouth (age range: 20 to 31, M=24; 7 self-identified as female and 1 as male).
+The system relies on engineered prompts such as splitting requests to make them faster, giving the model only a subset of the text to force localized edits, and giving the model the previous abstract representation of the story and the new one to help it understand the edits.
+See the appendix and/or the source code for more details.
+5. Study 1: Planning Using the Visualizations
+In our initial user study, we assessed how story visualizations aid the planning and reviewing phases of writing (Flower and Hayes, 1981).
+To ensure the study focused solely on visualization, the prototype only supported the
+features.
+We then compared this read-only version to a baseline text-only interface, asking participants to answer high-level planning and reviewing questions.
+Because no text had to be produced, we recruited participants with various creative writing experiences to reflect the many potential target users of the tool
+5.1. Method
+5.1.1. Participants
+We recruited 12 participants with various writing experiences from our social network and a mailing list of writers (age range: 20 to 37, M=25; 7 self-identified as male, 4 as female, and 1 preferred not to disclose).
+On a 5-point Likert scale ranging from 1-never to 5-always, they reported how often they write (Mdn=4), how often they use visuals while writing (Mdn=2), and how often they find visuals to help when writing (Mdn=3).
+For visual aids, some mentioned using timelines (N=5) and diagrams (N=4).
+They also reported different writing contexts, including academic (N=9), professional (N=7), and creative (N=6) writing.
+5.1.2. Apparatus
+Participants joined the study remotely via the Zoom conference software and accessed the experiment website using their own computers.
+The sessions lasted 1 hour and were audio and screen-recorded.
+Additionally, interactions with the experiment website were also recorded (e.g., features used, commands executed, and answers to questions).
+Participants received a CAD$20 gift card for their time.
+Our institutional ethics board approved the study.
+5.1.3. Tasks
+Tasks simulated what writers ask themselves when reviewing their craft.
+We extracted questions from best practices on revising fiction (Madden, 1988; Norton, 2011) and then revised them with a professional creative writing instructor.
+This led to two open-ended questions per story aspect: characters (“Could characters be combined without changing the outcome of the story?” and “Is any character too passive?”); locations (“Are there any locations which could be removed?” and “Are there moments where the spatial logic is broken or unclear?”); time and rhythm (“Is there a large gap between two actions that make the story progress?” and “Is there a scene you think could take place later/earlier in the story?”); and focalization (“If told from another character’s perspective, how would the story change?” and “What does the main character mention that no other character would?”).
+5.1.4. Stories
+We picked two short stories from the Tell Me A Story dataset (Huot et al., 2025), which consists of human-written stories.
+Stories were screened to be of similar length (~700 words) and featuring more than two speaking characters and location shifts.
+5.1.5. Procedure
+After signing the consent form and completing a demographic questionnaire, participants started with one of two conditions: visual story-writing or text-only.
+For the visual story-writing condition, they watched a one-minute tutorial and then tried the tool.
+For both conditions, they read the story fully before starting tasks.
+Participants had to think and answer questions aloud within five minutes, and then explain how they came up with the answer.
+After each condition, they completed a raw NASA TLX.
+At the end, they engaged in a semi-structured interview to share impressions, preferences, and strategies.
+Tasks, stories, and conditions were counterbalanced across participants.
+5.2. Results
+Participants were able to complete the tasks without difficulties in both conditions.
+However, how they accomplished the task varied.
+Themes are based on their relevance to our research questions, as per an interpretivist approach to reflexive thematic analysis (Braun and Clarke, 2022).
+Visualizations helped confirm, analyze, and explore.
+All participants mentioned using the visualizations to confirm their intuitions.
+Visualizations gave the “reassurance that my hunch is correct, and I’m not missing anything when answering the questions.” (P6) as they felt “it was mostly a confidence thing where I felt more sure in what I was saying” (P2).
+The visuals also helped “have conversations [about] deeper analysis […] and narrative structures” (P4).
+For example, participants used visuals to “remove this location simply because nobody was there” (P2), “identify who can know about each other” (P5),“figure out if [an event] was playing a part in the story” (P6), and see “how the entities interact with each other” (P12).
+P8 added it could be particularly useful “if you were editing or doing some kind of work on a piece of text where you’re like […] has everyone been given as much time to speak as I want them to.”
+Sometimes, visuals led to discoveries and reflections.
+For example, P11 mentioned “it kind of dawned on me by looking at the chronology that […] we actually don’t really get [character]’s motivations”.
+Similarly, P6 used the location view to “reflect on if [location] were to be changed to a different location, would that really have a significant impact on the story?
+And I realized that’s probably no.” P7 was “paging through the timeline and seeing how the characters jumped around and kind of weren’t even represented […] it actually showed me […] that it’s like, oh, a lot of moments in this story are sort of disconnected from one another”.
+P3 added “it does make me curious […] how would a story that features less locations look like”.
+Visual-driven search vs.
+skimming.
+Participants praised the visuals for their help finding passages.
+With the visuals, “it was way more tidier, similar kind of information were in similar places […] or visually looked similar” (P5) and “if I just click [an entity], it’ll just highlight the parts where this entity is being talked about […] I don’t have to look at other useless portions” (P12).
+In particular, participants contrasted this to their workflow with text, where “if I just remember, like vaguely, […] then I can probably pattern-match to the visualization faster than trying to figure out like, what is […] the keyword I could search to find the sequence of text.” (P3).
+In the text condition, participants mentioned how they “tried to hunt for the dialogues […] where there were quotations, and tried to see how many of them were said by [character]” (P4), and were finding locations by “looking for capitalized nouns […] so there’s always the chance that there was something else […] like, oh, the coast!” (P7).
+Impact on cognitive load.
+Results were inconclusive.
+Wilcoxon signed-rank did not find a significant difference for mental demand (p=.21), physical demand (p=1), temporal demand (p=.49), performance (p=.38), effort (p=.13), nor frustration (p=.16).
+Participants’ comments were also mixed.
+Some mentioned how the visualization reduced their perceived cognitive load because “[with text] I was trying to visualize the things in my head.
+But with [visuals], it kind of takes away that mental strain” (P12) and “it has all of the characters and locations listed out.
+If I were to use command F, I would have to put in whatever I can still remember, which oftentimes might not be the complete set”.
+Others appreciated being able to focus on other aspects because “the visualization […] lets you offload some of the more concrete stuff so that you could explicitly focus on things like character motivations, whereas I found, for the text-based one, I was constantly reskimming in order to refresh this mental schema” (P11).
+However, others mentioned that the system added to their cognitive load because “trying to learn [the tool] […] almost added this extra layer where I wasn’t engaging with the text” (P7).
+Mental model matches and mismatches.
+The visuals made sense to some participants because “I feel like my brain does this, […] but maybe not this clearly”.
+In fact, P9, who had the text condition after the visual condition, mentioned how they “laid down a timeline in [their] head, something akin to what we’ve seen in [the visual condition]” to accomplish the tasks.
+Yet, sometimes, there was a clear model mismatch.
+P8 was the most critical, explaining “I’m sure there’s people who just think like this, and this would be a really useful tool for them, because it is an extension of the way they process information.
+But for me, it’s like really not”.
+In fact, P8 mentioned often using visuals when writing but explained their process was different: “I will write those scenes on sticky notes and then I can start to arrange them in an order that gives me a plot […], and then I sometimes will use […] picture boards […] to remind me of a visual idea or concept”.
+Some other times, the mismatch happened at the semantic level: “calling it an event for me makes it feel like there’s a plot line that’s happening.
+But it wasn’t really that.” (P2) and “the timeline is very fine grained […] I wasn’t looking at fine-grained scenes […] but just in terms of more large-scale scenes” (P3).
+Complementarity of text and visuals Participants viewed the visualization as complementing the text because “you still need the text, the words, to fully appreciate the story” (P1).
+In fact, when asked how they did the task, participants reported using the views for 74% of tasks, their memory for 46%, and reading for 43% (adds up to more than 100% because two or more ways could be used).
+One reason is that the visuals pointed to text that then had to be verified.
+P4 explained “I was using text for the verification, because there were cases when the visualization was capturing something that could be potentially misleading”.
+Another reason is that the visualization could not capture everything.
+In particular, the visuals only represented explicit information: “It is really helpful in terms of managing kind of explicit things, like characters, entities, locations […] But it doesn’t quite capture things like motivations” (P11).
+6. Study 2: Editing and Free-form Writing
+The vision of visual story-writing is to complement the existing writing workflow.
+As visuals have shown to be helpful for reflecting on the text, we further explore how visual story-writing can assist writers during the “translating” and revision processes of the cognitive model of writing (Flower and Hayes, 1981).
+To explore this, we recruited experienced creative writers so that they could give us insights on the possible impact the tool would have on their writing workflows.
+This exploratory study used the fully-featured prototype and comprised two parts: the first part isolates the use of the three different views in a controlled setting, and the second part is a free-form creative writing task.
+We were interested in the following questions: Are the visual representations understood and helpful?
+Can people express their editing intents using visual story-writing?
+What is the impact of visual story-writing on exploration and creativity?
+6.1. Method
+6.1.1. Participants
+We recruited 8 creative writers from our social network and word-of-mouth (age range: 20 to 31, M=24; 7 self-identified as female and 1 as male).
+None participated in the first study.
 All had several years of experience with creative writing, including personal writing (N=7), fiction (N=4), scriptwriting (N=2), and poetry (N=2).
-On a scale from 1-never to 5-always, they reported how often they use visuals (e.g., mind maps, timelines) during their writing process (Mdn=2), how often they find visuals to help during the writing process (Mdn=3), and how frequently they use AI services to help during their writing activities (Mdn=1).
-Participants reported using a range of types of visuals to help their writing, including mood boards (N=4), timelines (N=3), mind maps (N=2), and reference imagery (N=2).
-5.2. Apparatus
-Participants joined the study remotely via the Zoom conference software.
-They accessed the experiment website using their own computer.
-The sessions were audio and screen recorded, and interactions with the experiment website were also recorded (e.g., features used, commands executed, and answers to questions).
-The sessions were about 45 minutes long.
-In appreciation for their time, participants received a CAD$15 gift card.
-The study was approved by our institutional ethics board.
-5.3. Design and Procedure
-Introduction (~2min)
-After being greeted, the participant was asked to complete the consent form and demographics questionnaire. They were then told that they would use a system to help them edit stories.
-Video Tutorial ((1 + 3 tasks) 30 seconds)
-Before each task, the participant watched a short video demonstrating the features they were about to use in the task.
-The video used a different text and different tasks to make sure the participant could not just reproduce what happened in the tutorial.
-Before the first task, the participant watched an additional video presenting the concept of the tool and the organization of the interface.
-Story Editing Tasks (3 tasks 5min)
-Immediately after watching the video tutorial, the participant were prompted to complete the task.
-We had three task blocks for each of the visual representation views: (1) entities and actions view tasks required removing an entity and then adding a new entity and an action with the main character; (2) location and entities view tasks required moving an entity to an existing location and then creating a new location and move another entity to it; (3) timeline view tasks required the reordering of events.
-The instructions for each task were phrased as questions and in a way that forced participants to engage with the story.
-For example, one of the tasks related to the timeline read “What if Jack loses his hat when he flies into the sky?”, forcing participants to understand enough about the story to know where those events occur and move the action to the right location.
-The tasks were counter-balanced across participants.
-Post-Task Questionnaire (3 tasks 2min)
-After each task, the participant first indicated how successful they felt in accomplishing the task on a 5-point semantic differential scale.
-Then, they answered five questions on a 5-point scale related to usability, understandability, and integration into their workflow.
-Free-form Exploration (5-15min)
-After accomplishing all the tasks and having a good grasp of the tool, the participant engaged in a free-form exploration.
-We provided the beginning of a story and asked the participant to continue the story with no instructions other than to “explore what could happen next”.
-Post-Study Questionnaire (2min)
-After completing the free-form part of the study, the participant answered one last questionnaire, which included the creativity support index (CSI (Cherry and Latulipe, 2014)) and the raw NASA TLX.
-Semi-structured Interview (10min)
-At the end, the participant engaged in a semi-structured interview with the experimenter.
-After giving their immediate impressions of the tool, they were asked about their experience with respect to the main categories of the CSI, like enjoyment, exploration, expressiveness, immersion, and satisfaction given the effort.
-They were also asked about each representation and their perceived overall advantages and disadvantages of the system.
+On a scale from 1-never to 5-always, they reported how often they use visuals (e.g., mind maps, timelines) (Mdn=2), how often they find visuals to help (Mdn=3), and how frequently they use AI services (Mdn=1) during writing.
+They reported prior use of visuals to help their writing, including mood boards (N=4), timelines (N=3), mind maps (N=2), and reference imagery (N=2).
+Participants are labelled W1 to W8 in the rest of the paper.
+6.1.2. Apparatus
+Same as 5.1.2 except the sessions were about 45 minutes long and participants received a CAD$15 gift card.
+6.1.3. Tasks
+The first part had three task blocks for each view: (1) entities and actions view tasks required removing an entity then adding a new entity and action involving the main character; (2) location and entities view tasks required moving an entity to an existing location, creating a new location, and move another entity to it; (3) timeline view tasks required reordering events.
+Instructions were phrased as questions, forcing participants to engage with the story.
+For example, a timeline task asked “What if Jack loses his hat when he flies into the sky?”.
+This required participants to understand the story’s events and move the action to the right location.
+Tasks were counterbalanced across participants.
+The second part involved a single writing task using a writing prompt.
+6.1.4. Procedure
+After completing the consent form and demographics questionnaire, participants watched a 30-second video introducing the tool.
+Before a task block, they watched a short video demonstrating the relevant features.
+After a block, participants rated their perceived success on a 5-point semantic differential scale and answered five questions related to usability, understandability, and workflow integration.
+Following all tasks, participants engaged in a free-form exploration to continue a story with no instructions other than to “explore what could happen next.” At the end, participants completed an exit questionnaire (including the creativity support index (CSI (Cherry and Latulipe, 2014)) and the raw NASA TLX) and engaged in a semi-structured interview to gather their experience and perceived advantages and disadvantages of the system.
 6.2. Results
-Participants completed the task and the free-form exploration without major difficulties.
+Participants completed the tasks and the free-form exploration without major difficulties.
 Below, we summarize the quantitative findings using descriptive statistics.
-Following previous recommendations for study designs similar to ours (Zhu and Kolassa, 2018; Masson et al., 2023), confidence intervals are calculated using the studentized bootstrapping method using 10,000 repetitions.
-Success rate
-On a 5-point semantic differential scale from 1-unsuccessful to 5-successful, participants rated their success M=4.25 (95% CI: [3.8, 4.6]) at accomplishing the tasks.
-Specifically, they were most successful at the task about adding and removing entities (M=4.5, 95% CI: [3.8, 5]), followed by changing the locations of entities (M=4.25, 95% CI: [3.4, 4.8]), and reordering events (M=4, 95% CI: [2.8, 4.7]).
-Creativity Support Index (CSI)
-Participants generated an average CSI score of 71.5 (SD=21.12) for the system and the tasks.
-Table 1 shows the breakdown of the CSI results.
-Enjoyment was both the most highly rated factor and also the factor considered most important by participants.
-Immersion was ranked second most important yet received the lowest score.
-Possible reasons for these ratings are discussed in Section 6.
+Following recommendations for similar study designs (Zhu and Kolassa, 2018; Masson et al., 2023), confidence intervals use the studentized bootstrap with 10,000 repetitions.
+Themes are based on their relevance to our research questions, as per an interpretivist approach to reflexive thematic analysis (Braun and Clarke, 2022).
 | Scale | Avg. Score (SD) | Avg. Counts (SD) |
@@
 | Overall CSI Score | 71.50 (21.12) |
-Workload (Raw NASA-TLX)
-Participants reported high performance (M=7.49 SD=1.95) for low mental demand (M=4.11, SD=2.38), effort (M=3.39, SD=2.17), frustration (M=2.5, SD=1.2), temporal demand (M=1.18, SD=1.04), and physical demand (M=1.18, SD=1.04).
-Questionnaire
-Overall, answers to the 5-point scale statements from 1-strongly disagree to 5-strongly agree were mostly positive, see Figure 5 for a breakdown of the results.
-Participants found the timeline the easiest view to understand (Mdn=4) and use (Mdn=4).
-Participants mostly agreed that the views helped accomplish the tasks (Mdn=5), that they would be useful within their writing process (Mdn=4), and that they would easily be integrated into their writing process (Mdn=4).
+Finding scenes and keeping track of events, entities, and locations.
+As a search tool, the results of this study corroborated those of the first study.
+Most participants agreed that “The visual representation was easy to understand” for all three views (Figure 7).
+W5 appreciated that you can “click on a person, click on a location, and then see stuff pop up” instead of “having to look through […] many paragraphs […] or doing keyword search”.
+W4 explained “Just having a visual representation helps, you know, just with the different perspectives”.
+W3 added “Moving around the timeline was a good way of, kind of figuring out different events”.
+W1 and W5 also mentioned how they usually struggle to keep track of everything while writing their stories and that the system could help them.
+For instance, W5 details that currently they “have booklets everywhere, and [they] have it annotated weird […] it’s a mess and this is so much easier.
+You got like, you know, your locations, your people […] keeping track of things.
+Organizing who’s where, what’s going on, being able to scroll through, like highlighting part of the text shown on the timeline […] making it easy to understand”.
+This was also reflected in participants reporting low mental demand on the NASA-TLX (M=7.49 SD=1.95).
+Specifying temporal, spatial, and entity-related edits.
+On a 5-point semantic differential scale from 1-unsuccessful to 5-successful, participants rated their success M=4.25 (95% CI: [3.8, 4.6]) at accomplishing the tasks.
+They were most successful at the task about adding and removing entities (M=4.5, 95% CI: [3.8, 5]), changing the locations of entities (M=4.25, 95% CI: [3.4, 4.8]), and reordering events (M=4, 95% CI: [2.8, 4.7]).
+Participants used an average of 2.42 visual edits to accomplish a task.
+In the free-form part, the most common visual edits were adding actions (36%), editing actions (26%), and editing traits (19%).
+Specifically, participants appreciated that the edits produced text as a good starting point.
+W2 mentioned “If I add in an extra character, it gives you just enough text to build that story with that new character”.
+W4 appreciated that it fixed inconsistencies when editing locations: “the fact that it also updated the sentence where the hat flew off his head that is extremely useful because that’s the sort of inconsistencies that can easily creep in when you’re making modifications like that”.
+Similarly, W6 expressed, “I really liked the ability to create relations between characters”.
+On the NASA-TLX, they also reported high-performance (M=7.49 SD=1.95) for low effort (M=3.39, SD=2.17), frustration (M=2.5, SD=1.2), temporal demand (M=1.18, SD=1.04), and physical demand (M=1.18, SD=1.04).
+For example, W6 further expressed that “[the tool] could be useful as a touch-up tool for when I want to make larger sweeping changes to a very long existing story […] I can confirm whether or not I want to keep each of these edits” stating that the system “made it a lot easier than any sort of other systems that I’ve used”.
+The idea of reviewing the changes was important for participants, as W3 explained “You’d still have to review it, but it made enough changes that it [the story] was able to flow well.”
+Helping explore and be creative.
+Participants generated an average CSI score of 71.5 (SD=21.12) (Table 2).
+They felt the tool was useful as a creative assistant, even when the text produced was not up to their standards.
+W6 explained “A lot of times, I am interested in how making a change to a given detail in a story would look […] I was appreciating [the tool] as sort of an exploratory tool where even if it doesn’t end up producing something that I would personally publish, I think it’s really good to give an example that I can then engage with”.
+Similarly, W5 mentioned “You can just like, throw in a verb or the general action, interaction between them and it was kind of spitballing ideas”.
+Regarding the timeline, W7 felt that “it’s an easy way to explore your options of what it sounds like if something happens here or if it sounds better when it’s there.”.
+They also mentioned how just looking at the visualization could help, for instance: “In terms of ideas, I could see how looking at the map, something could pop up into your head […] once you have, like, those objects laid out on a map, then that can help you spark inspiration”.
+Impact on the writing workflow.
+Perhaps due to the study setting, all participants heavily used the system during the free-form part.
+Most participants used the system to explore, “just to see what would happen” (W2), then read the suggested changes and kept them or undid them either by editing an event (W2, W6), undoing using the history graph (W3), deleting actions and characters (W1, W4, W6), or rewriting passages manually (W4, W8).
+When exploring, participants mostly used the actions and the location views.
+The timeline was only used to select specific scenes to edit.
+Participants also had different strategies when creating events, for example, W1, W4, and W7 wrote longer and detailed actions such as “Interrupts conversations, she says that she saw it too, and she is also confused” (W1) or explicitly writing a dialogue “Says ‘what are you doing here?”’ (W7), whereas others wrote one or two-word actions or descriptions.
+W8 had the most unique workflow in that they used the visuals to edit the text only once.
+Instead, they spent most of the session writing their story manually and would periodically refresh the visualization based on what they wrote and then reflect.
+They explained “In my most focused state I would just be focused on writing, but then when I was done, I was out of it, I would scan back and forth between text and the [visuals]”.
+Mismatches between the system and participants’ existing workflows.
+Questions related to integrating the system in the writing process were rated the lowest (but still with an Mdn=4).
+Similar to the findings from the first study, the most critical participants explained that the visuals did not match their workflow and the visualizations they already used.
+W4 mentioned “I really really like having a visual representation of the story, but this is not the visual representation I would choose […] I would like a combination of those two things [timeline and entity view] where those interlinks and connections appear within the timeline.”444While this visualization was not implemented in our system, it can be modelled with our framework: start with characters, connect by events, position by locations, and unfold by time.
+Similarly, W6 explained how a single timeline would not be enough for them: “my brain interprets timelines in stories where I enjoy having things happening in multiple places at the same time and having a single linear approach makes it very hard to have that sort of interesting kind of development”555Again, this can be modelled by our framework by unfolding by both time and temporality to visualize nonlinear narratives.
+Other than visual representations, W4 also mentioned they “don’t want the system overwriting […] because then it’s not me expressing myself through text” and instead, they would prefer “just highlighting the places where things need to change, or the things are inconsistent”.
+Participants would have liked more control in using the visual representations to edit the story.
+While participants could express their intents with the tool, there are nuances in how an idea can be expressed in terms of writing style and voice, which is not currently supported by our system.
+For example, when discussing adding actions by connecting two entities, W8 mentioned “I feel like that’s very limiting how you wanted to express what you want versus just typing it out in a sentence because it kind of format it in a specific manner”.
+Similarly, W7 explained “in terms of expressing myself, I would say that it does help […] but I don’t think it allows for like total freedom of expression […] if you had something, in your mind of how you wanted to say it, it might not come out that way”.
 7. Discussion
-We first discuss our results in light of participants’ comments before opening the discussion to other themes stemming from our observations and discussions with participants.
-Visual writing can help find scenes and keep track of events, entities, and locations
-Most participants agreed that “The visual representation was easy to understand” for all three views.
-Particularly, they mentioned how the views helped them.
-P5 appreciated that you can “click on a person, click on a location, and then see stuff pop up” instead of “having to look through […] many paragraphs […] or doing keyword search”.
-P4 explained “Just having a visual representation helps, you know, just with the different perspectives”.
-P3 added “Moving around the timeline was a good way of, kind of figuring out different events”.
-P1 and P5 also mentioned how they struggle to keep track of everything while writing their stories and that the system would help them keep track of that.
-For instance, P5 details that currently they “have booklets everywhere, and [they] have it annotated weird […] it’s a mess and this is so much easier.
-You got like, you know, your locations, your people […] keeping track of things.
-Organizing who’s where, what’s going on, being able to scroll through, like highlighting part of the text shown on the timeline […], making it easy to understand”.
-Visual writing can help specify temporal, spatial, and entity-related edits
-Participants reported high success rates when accomplishing the tasks and provided high ratings for statements such as “The visual representation helped me accomplish the tasks” and “The visual representation was easy to use”.
-This suggests that visual writing facilitates temporal, spatial, and entity-related edits through the views provided in our system.
-Participant’s comments further support these results.
-For example, P2 mentioned “If I add in an extra character, it gives you just enough text to build that story with that new character”.
-P4 appreciated that it made the story consistent when editing locations: “When I was dealing with the change where the hat was supposed to be left at home, the fact that it also updated the sentence where the hat flew off his head that is extremely useful because that’s the sort of inconsistencies that can easily creep in when you’re making modifications like that”.
-Similarly, P6 explained, “I really liked the ability to create relations between characters […] It’s something that made it a lot easier than any sort of other systems that I’ve used” and later added “[the tool] could be useful as a touch-up tool for when I want to make larger sweeping changes to a very long existing story […] I can confirm whether or not I want to keep each of these edits.”.
-This idea of reviewing the changes was important for participants, as P3 explained “You’d still have to review it, but it made enough changes that it [the story] was able to flow well.”
-Visual writing can help explore and be creative
-Many participants felt the tool was useful as a creative assistant.
-P6 explained “A lot of times, I am interested in how making a change to a given detail in a story would look […] I was appreciating it [the tool] as sort of an exploratory tool where even if it doesn’t end up producing something that I would personally publish, I think it’s really good to give an example that I can then engage with”.
-Similarly, P5 mentioned “You can just like, throw in a verb or the general action, interaction between them and it was kind of like spitballing ideas”.
-Regarding the timeline, P7 felt that “it’s an easy way to explore your options of what it sounds like if something happens here or if it sounds better when it’s there.”.
-They also mentioned how just looking at the visualization could help, for instance: “In terms of ideas, I could see how looking at the map, something could pop up into your head […] once you have, like, those objects laid out on a map, then that can help you spark inspiration”.
-6.1. Observations and Participants’ Comments
-6.1.1. Participants had different preferences
-For instance, P4 mentioned “I really really like having a visual representation of the story, but this is not the visual representation I would choose […] I would like a combination of those two things [timeline and entity view] where those interlinks and connections appear within the timeline.”.
-Similarly, P6 explained how a single timeline would not be enough for them: “my brain interprets timelines in stories where I enjoy having things happening in multiple places at the same time and having a single linear approach makes it very hard to have that sort of interesting kind of development”.
-Other than visual representations, P4 also mentioned they “don’t want the system overwriting […] because then it’s not me expressing myself through text” and instead, they would prefer “just highlighting the places where things need to change, or the things are inconsistent”.
-This corroborates previous research on creative support tools needing to offer “many paths and many styles” to do tasks because users are unpredictable and have different preferences (Shneiderman et al., 2006).
-6.1.2. Participants would have liked more control in using the visual representations to edit the story
-While participants could express their intents with the tool, there are nuances in how an idea can be expressed in terms of writing style and words used, which is not currently supported by our system.
-For example, when discussing adding actions by connecting two entities, P8 mentioned “I feel like that’s very limiting how you wanted to express what you want versus just typing it out in a sentence because it kind of format it in a specific manner”.
-Similarly, P7 explained “in terms of expressing myself, I would say that it does help […] but I don’t think it allows for like total freedom of expression […] if you had something like, in your mind of how you wanted to say it, it might not come out that way”.
-6.2. Limitations
-6.2.1. The participants in our study were experienced and hesitant to using AI
-All the participants in our study reported having multiple years of creative writing experience.
-Similarly, on a scale from 1-never to 5-always, participants rated their frequency of use of AI services a median of one, with five participants reporting never using AI services.
-P4 explained “I write because I want to, so if I want to be writing then there’s ideas that I have, and I want to express them.
-I don’t want a system to tell me what it is that I am going to be expressing”.
-This confirms previous findings that writers have different opinions towards AI (Hertzmann, 2018; Biermann et al., 2022) and those opinions might have impacted our results.
-6.2.2. The system and the representations do not scale to long stories
-One main limitation of the representations used in the system is that they quickly become too crowded and unusable for long stories.
-Currently, the system solves this problem through filtering mechanisms when the cursor hovers over a sentence or a scene.
-This results in hiding the elements irrelevant to the passage currently being selected.
-But this only works for local edits.
-When the edits need to happen between scenes far apart in the story, then it becomes challenging to find the right scene or the correct location.
-We believe that, for these cases, new aggregated representations and semantic zooming would be necessary.
-6.2.3. The system might not have been robust enough to test the full potential of visual writing
-While the system was optimized to be robust, it still depends on a large language model to understand the story and the instructions properly.
-In our tests, but also during the study, we observed the system sometimes responding unexpectedly by modifying elements it should not or by refusing to do certain modifications.
-Additionally, despite our best efforts to parallelize requests and speed up the execution, some manipulations were still slow to execute (up to 10-15 seconds at times).
-These issues might have confounded the results of our study.
-For example, it is known that to promote user exploration, interactions should be fast and with effects immediately visible (Masson et al., 2022; Shneiderman et al., 2006; Shneiderman, 1983).
+We propose visual story-writing, an approach that uses visual representations of story elements to support writing and revising narrative texts.
+Despite how little writing support tools have used visualizations (Zhao et al., 2025), our findings show that they hold great promise to help review and ideate stories.
+The space of possibilities is large, and our framework of story constructs is an initial attempt at structuring it.
+In our study, we looked at a subset of this space, using simple representations that did not cover all story elements.
+However, we encourage others to continue this line of research.
+One question with tools like this is how it might change the role of writers.
+AI-assisted writing tools tend to make us write less and review and edit more (Buschek, 2024).
+We believe visual story-writing can accommodate both traditional and “editorial” writing roles: in its simplest form, visual story-writing is automatizing the creation of visualizations writers were already using (Zhao et al., 2025; Ackerman, 2017; Neuwirth and Kaufer, 1989; SaveTheCat, 2014; Articy, 2014), with the added bonus of serving as input, if the writer desires it.
+Another question is how visual story-writing could be integrated with other writing support tools.
+We view visual story-writing as complementary to tools that rely on textual prompts (e.g., ChatGPT or Grammarly).
+Like work on direct manipulation with LLMs has shown (Masson et al., 2024), we expect visual story-writing to be preferable when writers have a clearer idea of the edit they want to perform (i.e., “move this character there”) because it helps refer to story elements unambiguously.
+Whereas textual prompts are preferable when the goal is vague (e.g., “make this story sound better”).
+Finally, our work targets the creative writing of narratives that involve tangible elements like characters and locations.
+It applies not only to novels but also to screenwriting for films, TV shows, video games, and theatre plays.
+However, it is unclear how visual story-writing could work in domains like academic writing, where the elements manipulated are not so tangible.
 7.1. Future Work
-6.3.1. Visual writing to edit writing style and story plots
-Creative writing is more than events, actors, locations, and time.
-We believe that visual writing could also help editing other aspects, such as the style of the writing.
+7.1.1. Extending the framework and exploring new visual interactions
+Our framework (§3) operates at a different conceptual level than typical visualization frameworks, such as the grammar of graphics (Wilkinson et al., 2005).
+It combines story elements into constructs that are then linked to visuals without prescribing how to extract story elements or what the visualizations should look like, allowing more freedom in implementation.
+For example, our system represents locations as bubbles instead of a traditional map.
+Alternatively, images could indicate what an object, character, or location looks like, allowing manipulation (e.g., changing a character’s hair colour) or emphasis on certain descriptive points.
+More work is needed to identify which constructs can best serve writers within visual story-writing workflows, and how to best represent them.
+For example, inter-character relationships are important in a story (Bochner et al., 1997) and a view constructed by connecting actors would help visualize and edit these relationships.
+Additionally, the framework can be extended to include story elements beyond those highlighted by Bal (Bal, 1997).
+For example, ”emotions” could be a basic building block, as used in StoryPrint (Watson et al., 2019), and ”motivations” could be another, as suggested by P11 during the first study.
+In fact, other narratology models, such as Propp’s morphology of the folk tale (Propp and Pirkova-Jakobson, 2010), Todorov’s principles of narratives (Todorov, 1971), or the contemporary concepts on narrative time by Hume (Hume, 2014) would lead to a different framework.
+7.1.2. Visual editing of writing style and story plots
+Creative writing encompasses more than just events, actors, locations, and time.
+The approach could also help edit other aspects, such as writing style.
 As proposed by Kempen et al.
-(1987), the syntax tree of the sentences could be made visible and be manipulable to restructure sentences.
-Alternatively, the tone, the length of the sentences, or the sentence structure could be represented using histograms or by embellishing the text.
-Then, manipulating these visuals would allow tweaking the style of the story.
-At a higher level, visual writing could also help revise the plot of the story.
-Similar to work on manipulating outlines (Dang et al., 2022a; Zhang et al., 2023), the idea would be to represent the main plot events of the story and then make them manipulable to reorder or change them.
-6.3.2. Exploring new visual interactions
-The idea of visual writing is broader than the representations and interactions proposed by our system.
-Any visual representation better suited for a task could be beneficial.
-For example, one could imagine using images to give the system an indication of what an object, character, or location look like.
-Then, the image could be modified to update the story (e.g., painting the hair of a character blond would edit the story so that the description matches the new hair colour).
-The system could also allow highlighting part of the image to indicate which aspect should be more or less emphasized in the description.
-Similarly, while our implementation uses LLM, visual writing could be implemented without and made more or less autonomous (Morris et al., 2024).
-Using LLM comes about with many challenges related to biases, privacy, and plagiarism (Weidinger et al., 2021; Meyer et al., 2023).
-An alternative, suggested by P4, could be to have the visual writing system only highlight the parts that should be modified to help the writer identify and edit them themselves.
+(1987), manipulating the syntax tree can restructure sentences.
+Alternatively, tone, sentence length, or structure could be represented with histograms or text embellishments, allowing style adjustments through visual manipulation.
+At a higher level, visual story-writing can aid revising the plot.
+Similar to work on manipulating outlines (Dang et al., 2022; Zhang et al., 2023), main plot events can be visually represented to be reordered or changed.
 7.1.3. Suggestions and creativity support in the visual space
 The mental processes involved in interpreting visuals differ from those involved in processing language (Paivio, 1990).
-Future work could explore the impact on creativity of writing stories visually rather than with text.
-Similarly, traditional creativity support techniques for writers would be interesting to adapt to the visual space.
+Future work could explore the impact on the creativity of writing stories through visuals rather than with text.
+Similarly, typical writing support features would be interesting to adapt to the visual space.
 For instance, there are many benefits of phrase suggestions (Buschek et al., 2021; Bhat et al., 2023), and future work could investigate if there are similar benefits to suggestions in the visual space.
-These visual suggestions would be shown when an interaction with a visual element is initiated to highlight different alternatives and to “auto-complete” the interaction.
+These visual suggestions would essentially try to “auto-complete” an interaction initiated on a visual element.
+7.1.4. Supporting long stories
+For long stories, the visuals have to find the right balance between overview and details.
+Our prototype attempts to show the details of all the events and entities and rely on interactions to declutter the views, by merging overlapping actions and allowing users to hover or select a passage to filter what is shown.
+If details are prioritized, this view could be unfolded by temporality to “spread out” the interactions over time.
+Unfolding again by time would visualize complex nonlinear narratives.
+Alternatively, interaction techniques such as details-on-hover or semantic zooming could also help manage longer stories.
+7.1.5. Visualization-builder to support custom visuals
+The studies showed creative writers have various preferences in terms of visuals.
+One way to support these workflows would be to help users construct their own visuals.
+Our procedural framework could inform such a view-builder: the eight story elements would be the basic building blocks, and users could construct a view by applying the different operators.
+Once the view is constructed, it could become interactive and automatically populated from the text.
+7.2. Limitations
+7.2.1. Participants’ experience and reservation towards AI might have impacted our results
+In study 1, half the participants had no creative writing experience, meaning they were possibly at a disadvantage to accomplish the tasks.
+Conversely, all participants in study 2 had years of creative experience, and, while they could comment on how visual story-writing could be integrated in their workflows, the results could not tell us about the benefits of the tool for beginners.
+Similarly, some participants expressed reservations towards AI and reported rarely, if ever, using it.
+W4 explained “I write because I want to, so if I want to be writing then there’s ideas that I have, and I want to express them.
+I don’t want a system to tell me what it is that I am going to be expressing”.
+Considering writers’ have strong opinions about AI (Hertzmann, 2018; Biermann et al., 2022), this might have biased the results.
+7.2.2. The system might not have been robust enough to test the full potential of visual story-writing
+The system relied on an LLM that occasionally responded unexpectedly, such as modifying unintended elements or refusing to do certain modifications.
+Additionally, some manipulations were slow to execute (up to 10-15 seconds) despite our efforts to parallelize requests.
+These issues may have affected our results as interactions should be seamless and paired with immediate effects to promote user exploration (Masson et al., 2022; Shneiderman et al., 2006; Shneiderman, 1983).
 8. Conclusion
-We proposed visual writing, an approach to writing by manipulating visual representations to express editing intents.
-Through a taxonomy of possible visual representations based on narratology and a prototype implementation of some of these visual representations, we demonstrated one possible design for visual writing tools.
-A study with eight creative writers showed the potential of creative writing to help keep track of story elements, rapidly specify edits, and explore story variations in a way that encourages creativity.
-Broadly, our work advocates for a new generation of writing support tools that allow text editing through direct manipulation of visual representations.
+We proposed visual story-writing, an approach to support writing by reviewing and manipulating visual representations.
+In doing so, we defined a framework to help inform the design of story visualizations.
+By applying this framework, we implemented a prototype system demonstrating one possible design for visual story-writing tools.
+Two studies covering the different aspects of the writing process showed the potential of creative writing to help keep track of story elements, rapidly specify edits, and explore story variations in a way that encourages creativity.
+Broadly, our work advocates for a new generation of writing support tools that embed visualization to help review and edit text narratives.
 Acknowledgements.
-This work was supported by NSERC under grant RGPIN-2018-05072.References
+We thank Daniel Aureliano Newman for his help revising the questions from study 1.
+This project was undertaken thanks to funding from IVADO, the Canada First Research Excellence Fund, and NSERC Discovery Grant RGPIN-2018-05072.References
 - (1)
 - Ackerman (2017) Angela Ackerman. 2017. The Efficient Writer: Using Timelines to Organize Story Details.
-- Arnold et al. (2021) Kenneth C Arnold, April M Volzer, and Noah G Madrid. 2021. Generative Models Can Help Writers without Writing for Them. In IUI Workshops.
+- Arnold et al.
+(2021) Kenneth C Arnold, April M Volzer, and Noah G Madrid.
+2021.
+Generative Models Can Help Writers without Writing for Them.
+In Joint Proceedings of the ACM IUI 2021 Workshops.
+CEUR, College Station, USA, 8.
+- Articy (2014) Articy Software GmbH & Co 2014. Articy: narrative design for interactive projects. Articy Software GmbH & Co. https://www.articy.com
 - Bal (1997) Mieke Bal. 1997. Narratology: Introduction to the Theory of Narrative (2nd edition ed.). University of Toronto Press, Scholarly Publishing Division, Toronto.
+- Bal (2021) Mieke Bal. 2021. Narratology in Practice. University of Toronto Press, Toronto.
+- Bensaid et al.
+(2021) Eden Bensaid, Mauro Martino, Benjamin Hoover, and Hendrik Strobelt.
+2021.
+FairyTailor: A Multimodal Generative Framework for Storytelling.
+arXiv:2108.04324 [cs] doi:10.48550/arXiv.2108.04324
 - Bhat et al.
@@
 Association for Computing Machinery, New York, NY, USA, 436–452.
-https://doi.org/10.1145/3581641.3584060
+doi:10.1145/3581641.3584060
 - Biermann et al.
@@
 ACM, Virtual Event Australia, 1209–1227.
-https://doi.org/10.1145/3532106.3533506
+doi:10.1145/3532106.3533506
+- Bochner et al.
+(1997) Arthur P.
+Bochner, Carolyn Ellis, and Lisa M.
+Tillman-Healy.
+1997.
+Relationships as Stories.
+In Handbook of Personal Relationships: Theory, Research and Interventions, 2nd Ed.
+John Wiley & Sons, Inc., Hoboken, NJ, US, 307–324.
 - Brath (2021) Richard Brath. 2021. Surveying Wonderland for Many More Literature Visualization Techniques. https://arxiv.org/abs/2110.08584v1.
-- Burnett and McIntyre (1995) Margaret M. Burnett and David W. McIntyre. 1995. Visual Programming. Computer 28, 03 (March 1995), 14–16. https://doi.org/10.1109/MC.1995.10027
+- Braun and Clarke (2022) Virginia Braun and Victoria Clarke. 2022. Thematic Analysis: A Practical Guide. SAGE Publications Ltd, Los Angeles.
+- Bressa et al.
+(2024) Nathalie Bressa, Jordan Louis, Wesley Willett, and Samuel Huron.
+2024.
+Input Visualization: Collecting and Modifying Data with Visual Representations.
+In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24).
+Association for Computing Machinery, New York, NY, USA, Article 499, 18 pages.
+doi:10.1145/3613904.3642808
+- Burnett and McIntyre (1995) Margaret M. Burnett and David W. McIntyre. 1995. Visual Programming. Computer 28, 03 (March 1995), 14–16. doi:10.1109/MC.1995.10027
+- Buschek (2024) Daniel Buschek.
+2024.
+Collage Is the New Writing: Exploring the Fragmentation of Text and User Interfaces in AI Tools.
+In Proceedings of the 2024 ACM Designing Interactive Systems Conference (DIS ’24).
+Association for Computing Machinery, New York, NY, USA, 2719–2737.
+doi:10.1145/3643834.3660681
 - Buschek et al.
@@
 Association for Computing Machinery, New York, NY, USA, 1–13.
-https://doi.org/10.1145/3411764.3445372
+doi:10.1145/3411764.3445372
 - Calderwood et al.
@@
 How Novelists Use Generative Language Models: An Exploratory User Study..
-In HAI-GEN+ User2agent@ IUI.
+In IUI ’20 Workshops.
+CEUR, Cagliari, Italy, 5.
 - Carroll et al.
@@
 Association for Computing Machinery, New York, NY, USA, 127–136.
-https://doi.org/10.1145/1640233.1640255
-- Chakrabarty et al.
-(2024) Tuhin Chakrabarty, Vishakh Padmakumar, Faeze Brahman, and Smaranda Muresan.
-2024.
-Creativity Support in the Age of Large Language Models: An Empirical Study Involving Emerging Writers.
-arXiv:2309.12570 [cs]
+doi:10.1145/1640233.1640255
 - Cherry and Latulipe (2014) Erin Cherry and Celine Latulipe.
@@
 21, 4 (June 2014), 21:1–21:25.
-https://doi.org/10.1145/2617588
+doi:10.1145/2617588
 - Chung et al.
@@
 ACM, New Orleans LA USA, 1–19.
-https://doi.org/10.1145/3491102.3501819
+doi:10.1145/3491102.3501819
+- Chung et al.
+(2025) John Joon Young Chung, Melissa Roemmele, and Max Kreminski.
+2025.
+Toyteller: AI-powered Visual Storytelling Through Toy-Playing with Character Symbols.
+In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25).
+Association for Computing Machinery, New York, NY, USA, 1–23.
+doi:10.1145/3706598.3713435
+- Cinnamon (2017) Nadieh Bremer-Visual Cinnamon. 2017. All Fights in Dragon Ball Z. https://dragonballz.visualcinnamon.com.
 - Comberg ([n. d.]) David Comberg. [n. d.]. Kurt Vonnegut on Shape of Stories. https://www.youtube.com/watch?v=oP3c1h8v2ZQ
 - Dang et al.
-(2022a) Hai Dang, Karim Benharrak, Florian Lehmann, and Daniel Buschek.
-2022a.
+(2022) Hai Dang, Karim Benharrak, Florian Lehmann, and Daniel Buschek.
+2022.
 Beyond Text Generation: Supporting Writers with Continuous Automatic Text Summaries.
@@
 ACM, Bend OR USA, 1–13.
-https://doi.org/10.1145/3526113.3545672
+doi:10.1145/3526113.3545672
 - Dang et al.
@@
 Choice Over Control: How Users Write with Large Language Models Using Diegetic and Non-Diegetic Prompting.
-In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems.
-1–17.
-https://doi.org/10.1145/3544548.3580969 arXiv:2303.03199 [cs]
-- Dang et al.
-(2022b) Hai Dang, Lukas Mecke, Florian Lehmann, Sven Goller, and Daniel Buschek.
-2022b.
-How to Prompt?
-Opportunities and Challenges of Zero- and Few-Shot Learning for Human-AI Interaction in Creative Applications of Generative Models.
-https://arxiv.org/abs/2209.01390v1.
-- Freytag (1895) Gustav Freytag. 1895. Technique of the drama: An exposition of dramatic composition and art. S. Griggs.
+In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23).
+Association for Computing Machinery, New York, NY, USA, 1–17.
+doi:10.1145/3544548.3580969
+- Flower and Hayes (1981) Linda Flower and John R. Hayes. 1981. A Cognitive Process Theory of Writing. College Composition and Communication 32, 4 (1981), 365–387. arXiv:356600 doi:10.2307/356600
+- Freytag (2004) Gustav Freytag. 2004. Technique of the Drama: An Exposition of Dramatic Composition and Art. University Press of the Pacific.
+- Genette et al. (1990) Gérard Genette, Nitsa Ben-Ari, and Brian McHale. 1990. Fictional narrative, factual narrative. Poetics today 11, 4 (1990), 755–774.
 - Gobert and Beaudouin-Lafon (2023) Camille Gobert and Michel Beaudouin-Lafon.
@@
 Association for Computing Machinery, New York, NY, USA, 1–16.
-https://doi.org/10.1145/3586183.3606817
+doi:10.1145/3586183.3606817
 - Hempel et al.
@@
 Association for Computing Machinery, New York, NY, USA, 281–292.
-https://doi.org/10.1145/3332165.3347925
+doi:10.1145/3332165.3347925
 - Hertzmann (2018) Aaron Hertzmann. 2018. Can Computers Create Art? https://arxiv.org/abs/1801.04486v6.
@@
 ACM, Pittsburgh PA USA, 74–94.
-https://doi.org/10.1145/3563657.3596000
+doi:10.1145/3563657.3596000
+- Hsu et al.
+(2019a) Ting-Yao Hsu, Yen-Chia Hsu, and Ting-Hao (Kenneth) Huang.
+2019a.
+On How Users Edit Computer-Generated Visual Stories.
+In Extended Abstracts of the 2019 CHI Conference on Human Factors in Computing Systems (CHI EA ’19).
+Association for Computing Machinery, New York, NY, USA, 1–6.
+doi:10.1145/3290607.3312965
+- Hsu et al.
+(2019b) Ting-Yao Hsu, Chieh-Yang Huang, Yen-Chia Hsu, and Ting-Hao Huang.
+2019b.
+Visual Story Post-Editing.
+In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, Anna Korhonen, David Traum, and Lluís Màrquez (Eds.).
+Association for Computational Linguistics, Florence, Italy, 6581–6586.
+doi:10.18653/v1/P19-1658
+- Huang et al.
+(2016) Ting-Hao Kenneth Huang, Francis Ferraro, Nasrin Mostafazadeh, Ishan Misra, Aishwarya Agrawal, Jacob Devlin, Ross Girshick, Xiaodong He, Pushmeet Kohli, Dhruv Batra, C.
+Lawrence Zitnick, Devi Parikh, Lucy Vanderwende, Michel Galley, and Margaret Mitchell.
+2016.
+Visual Storytelling.
+In Proceedings of the 2016 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Kevin Knight, Ani Nenkova, and Owen Rambow (Eds.).
+Association for Computational Linguistics, San Diego, California, 1233–1239.
+doi:10.18653/v1/N16-1147
+- Hulstein et al.
+(2023) Golina Hulstein, Vanessa Peña-Araya, and Anastasia Bezerianos.
+2023.
+Geo-Storylines: Integrating Maps into Storyline Visualizations.
+IEEE Transactions on Visualization and Computer Graphics 29, 1 (Jan.
+2023), 994–1004.
+doi:10.1109/TVCG.2022.3209480
+- Hume (2014) Kathryn Hume. 2014. Fantasy and Mimesis: Responses to Reality in Western Literature. Routledge, Abingdon New York.
+- Huot et al.
+(2025) Fantine Huot, Reinald Kim Amplayo, Jennimaria Palomaki, Alice Shoshana Jakobovits, Elizabeth Clark, and Mirella Lapata.
+2025.
+Agents’ Room: Narrative Generation through Multi-step Collaboration.
+arXiv:2410.02603 [cs] doi:10.48550/arXiv.2410.02603
 - Hutchins et al.
@@
 1985), 311–338.
-https://doi.org/10.1207/s15327051hci0104_2
-- Kempen et al. (1987) Gerard Kempen, G. Ambeek, Peter Desain, L. Konst, and Koenraad De Smedt. 1987. Author Environments: Fifth Generation Text Processors. (Jan. 1987).
+doi:10.1207/s15327051hci0104_2
+- Kempen et al.
+(1987) Gerard Kempen, Gert Anbeek, Peter Desain, Leo Konst, and Koenraad De Smedt.
+1987.
+Author Environments: Fifth Generation Text Processors.
+In ESPRIT’86: Results and Achievements.
+Elsevier Science Publishers, Brussels, Belgium, 8.
 - Kim et al.
@@
 2018), 595–604.
-https://doi.org/10.1109/TVCG.2017.2744118
+doi:10.1109/TVCG.2017.2744118
 - Ko and Myers (2006) Amy J.
@@
 Association for Computing Machinery, New York, NY, USA, 387–396.
-https://doi.org/10.1145/1124772.1124831
+doi:10.1145/1124772.1124831
 - Lee et al.
@@
 A Design Space for Intelligent and Interactive Writing Assistants.
-https://doi.org/10.1145/3613904.3642697 arXiv:2403.14117 [cs]
+arXiv:2403.14117 [cs] doi:10.1145/3613904.3642697
+- Liveley (2021) Genevieve Liveley. 2021. Narratology. Oxford University Press, Oxford.
+- Madden (1988) David Madden. 1988. Revising Fiction: A Handbook for Writers. Plume, New York.
 - Maharjan et al.
@@
 Association for Computational Linguistics, New Orleans, Louisiana, 259–265.
-https://doi.org/10.18653/v1/N18-2042
-- Marshall (2014) Colin Marshall. 2014. How J.K. Rowling Plotted Harry Potter with a Hand-Drawn Spreadsheet | Open Culture.
+doi:10.18653/v1/N18-2042
 - Marti et al.
@@
 Association for Computing Machinery, New York, NY, USA, 509–519.
-https://doi.org/10.1145/3172944.3172972
+doi:10.1145/3172944.3172972
+- Masson et al.
+(2025) Damien Masson, Young-Ho Kim, and Fanny Chevalier.
+2025.
+Textoshop: Interactions Inspired by Drawing Software to Facilitate Text Editing.
+In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’25).
+Association for Computing Machinery, New York, NY, USA, Article 775, 14 pages.
+doi:10.1145/3706598.3713862
 - Masson et al.
@@
 Association for Computing Machinery, New York, NY, USA, 1–14.
-https://doi.org/10.1145/3586183.3606762
+doi:10.1145/3586183.3606762
 - Masson et al.
@@
 Association for Computing Machinery, New York, NY, USA, 1–16.
-https://doi.org/10.1145/3613904.3642462
+doi:10.1145/3613904.3642462
 - Masson et al.
@@
 Association for Computing Machinery, New York, NY, USA, 1–13.
-https://doi.org/10.1145/3491102.3501895
+doi:10.1145/3491102.3501895
 - Mayer et al.
@@
 2018), 127:1–127:28.
-https://doi.org/10.1145/3276497
-- Meyer et al.
-(2023) Jesse G.
-Meyer, Ryan J.
-Urbanowicz, Patrick C.
-N.
-Martin, Karen O’Connor, Ruowang Li, Pei-Chen Peng, Tiffani J.
-Bright, Nicholas Tatonetti, Kyoung Jae Won, Graciela Gonzalez-Hernandez, and Jason H.
-Moore.
-2023.
-ChatGPT and Large Language Models in Academia: Opportunities and Challenges.
-BioData Mining 16, 1 (July 2023), 20.
-https://doi.org/10.1186/s13040-023-00339-9
-- Morris et al.
-(2024) Meredith Ringel Morris, Jascha Sohl-dickstein, Noah Fiedel, Tris Warkentin, Allan Dafoe, Aleksandra Faust, Clement Farabet, and Shane Legg.
-2024.
-Levels of AGI for Operationalizing Progress on the Path to AGI.
-https://doi.org/10.48550/arXiv.2311.02462 arXiv:2311.02462 [cs]
+doi:10.1145/3276497
 - Munroe (2009) Randall Munroe. 2009. Movie Narrative Charts. https://xkcd.com/657/.
-- Myers (1986) B. A. Myers. 1986. Visual Programming, Programming by Example, and Program Visualization: A Taxonomy. SIGCHI Bull. 17, 4 (April 1986), 59–66. https://doi.org/10.1145/22339.22349
+- Myers (1986) B. A. Myers. 1986. Visual Programming, Programming by Example, and Program Visualization: A Taxonomy. SIGCHI Bull. 17, 4 (April 1986), 59–66. doi:10.1145/22339.22349
 - Neuwirth and Kaufer (1989) C.
@@
 ACM Press, Pittsburgh, Pennsylvania, United States, 319–341.
-https://doi.org/10.1145/74224.74250
+doi:10.1145/74224.74250
 - NextUI (2021) nextui-org 2021. NextUI. nextui-org. https://nextui.org/
+- North and Shneiderman (2001) Chris North and Ben Shneiderman. 2001. A Taxonomy of Multiple Window Coordinations. (2001), 8.
+- Norton (2011) Scott Norton. 2011. Developmental Editing: A Handbook for Freelancers, Authors, and Publishers (illustrated edition ed.). University of Chicago Press, Chicago London.
 - OpenAI (2023) OpenAI. 2023. OpenAI Node API Library. OpenAI.
-- Paivio (1990) Allan Paivio. 1990. Mental Representations: A Dual Coding Approach. Oxford University Press.
-- Pope (1998) Thomas Pope. 1998. Good scripts, bad scripts: learning the craft of screenwriting through 25 of the best and worst films in history. Three Rivers Press.
+- Paivio (1990) Allan Paivio. 1990. Mental Representations: A Dual Coding Approach. Oxford University Press, Oxford.
+- Pope (1998) Tom Pope. 1998. Good Scripts, Bad Scripts: Learning the Craft of Screenwriting Through 25 of the Best and Worst Films in Hi Story. Three Rivers Press, New York.
+- Propp and Pirkova-Jakobson (2010) V. Propp and Svatava Pirkova-Jakobson. 2010. Morphology of the Folk Tale. University of Texas Press.
 - Qiang and Bingjie (2016) Lu Qiang and Chai Bingjie.
@@
 IEEE, Chongqing, China, 211–218.
-https://doi.org/10.1109/CW.2016.43
+doi:10.1109/CW.2016.43
 - React (2013) Meta 2013. React – A JavaScript Library for Building User Interfaces. Meta. https://reactjs.org/
@@
 2016), 1–12.
-https://doi.org/10.1140/epjds/s13688-016-0093-1
+doi:10.1140/epjds/s13688-016-0093-1
 - Reza et al.
@@
 ABScribe: Rapid Exploration & Organization of Multiple Writing Variations in Human-AI Co-Writing Tasks Using Large Language Models.
-https://doi.org/10.1145/3613904.3641899 arXiv:2310.00117 [cs]
+arXiv:2310.00117 [cs] doi:10.1145/3613904.3641899
 - SaveTheCat (2014) Blake Snyder Enterprises 2014. Save The Cat! Story Structure Software. Blake Snyder Enterprises. https://savethecat.com/save-the-cat-software
-- Shneiderman (1983) Shneiderman. 1983. Direct Manipulation: A Step Beyond Programming Languages. Computer 16, 8 (Aug. 1983), 57–69. https://doi.org/10.1109/MC.1983.1654471
+- Shneiderman (1983) Shneiderman. 1983. Direct Manipulation: A Step Beyond Programming Languages. Computer 16, 8 (Aug. 1983), 57–69. doi:10.1109/MC.1983.1654471
 - Shneiderman (1982) Ben Shneiderman.
@@
 Behaviour & Information Technology 1, 3 (July 1982), 237–256.
-https://doi.org/10.1080/01449298208914450
+doi:10.1080/01449298208914450
 - Shneiderman et al.
@@
 International Journal of Human–Computer Interaction 20, 2 (May 2006), 61–77.
-https://doi.org/10.1207/s15327590ijhc2002_1
+doi:10.1207/s15327590ijhc2002_1
 - SlateJS (2016) Slate.js 2016. Slate.js. Slate.js. https://www.slatejs.org
@@
 Association for Computing Machinery, New York, NY, USA, Article 644, 26 pages.
-https://doi.org/10.1145/3613904.3642400
+doi:10.1145/3613904.3642400
 - Tanahashi and Kwan-Liu Ma (2012) Y.
@@
 2012), 2679–2688.
-https://doi.org/10.1109/TVCG.2012.212
+doi:10.1109/TVCG.2012.212
+- Todorov (1971) Tzvetan Todorov. 1971. The 2 Principles of Narrative. Diacritics 1, 1 (1971), 37–44. arXiv:464558 doi:10.2307/464558
+- Wang et al.
+(2018) Jing Wang, Jianlong Fu, Jinhui Tang, Zechao Li, and Tao Mei.
+2018.
+Show, Reward and Tell: Automatic Generation of Narrative Paragraph from Photo Stream by Adversarial Training.
+In Proceedings of the Thirty-Second AAAI Conference on Artificial Intelligence and Thirtieth Innovative Applications of Artificial Intelligence Conference and Eighth AAAI Symposium on Educational Advances in Artificial Intelligence (AAAI’18/IAAI’18/EAAI’18).
+AAAI Press, New Orleans, Louisiana, USA, 7396–7403.
+- Wang Baldonado et al.
+(2000) Michelle Q.
+Wang Baldonado, Allison Woodruff, and Allan Kuchinsky.
+2000.
+Guidelines for using multiple views in information visualization.
+In Proceedings of the working conference on Advanced visual interfaces.
+ACM, Palermo Italy, 110–119.
+doi:10.1145/345513.345271
 - Watson et al.
@@
 Association for Computing Machinery, New York, NY, USA, 303–311.
-https://doi.org/10.1145/3301275.3302302
-- Weidinger et al.
-(2021) Laura Weidinger, John Mellor, Maribeth Rauh, Conor Griffin, Jonathan Uesato, Po-Sen Huang, Myra Cheng, Mia Glaese, Borja Balle, Atoosa Kasirzadeh, Zac Kenton, Sasha Brown, Will Hawkins, Tom Stepleton, Courtney Biles, Abeba Birhane, Julia Haas, Laura Rimell, Lisa Anne Hendricks, William Isaac, Sean Legassick, Geoffrey Irving, and Iason Gabriel.
-2021.
-Ethical and Social Risks of Harm from Language Models.
-https://doi.org/10.48550/arXiv.2112.04359 arXiv:2112.04359 [cs]
-- Williams ([n. d.]) Eric Williams. [n. d.]. How to View and Appreciate Great Movies. https://www.thegreatcourses.com/courses/how-to-view-and-appreciate-great-movies
+doi:10.1145/3301275.3302302
+- Wilkinson et al. (2005) Leland Wilkinson, D. Wills, D. Rope, A. Norton, and R. Dubbs. 2005. The Grammar of Graphics (2nd edition ed.). Springer, New York.
+- Williams (2018) Eric R. Williams. 2018. How to View and Appreciate Great Movies.
+- Yan et al.
+(2023) Zihan Yan, Chunxu Yang, Qihao Liang, and Xiang ’Anthony’ Chen.
+2023.
+XCreation: A Graph-based Crossmodal Generative Creativity Support Tool.
+In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology.
+ACM, San Francisco CA USA, 1–15.
+doi:10.1145/3586183.3606826
 - Yuan et al.
@@
 Association for Computing Machinery, New York, NY, USA, 841–852.
-https://doi.org/10.1145/3490099.3511105
+doi:10.1145/3490099.3511105
+- Zamfirescu-Pereira et al.
+(2023) J.D.
+Zamfirescu-Pereira, Richmond Y.
+Wong, Bjoern Hartmann, and Qian Yang.
+2023.
+Why Johnny Can’t Prompt: How Non-AI Experts Try (and Fail) to Design LLM Prompts.
+In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23).
+Association for Computing Machinery, New York, NY, USA, 1–21.
+doi:10.1145/3544548.3581388
 - Zhang et al.
@@
 ACM, San Francisco CA USA, 1–30.
-https://doi.org/10.1145/3586183.3606800
+doi:10.1145/3586183.3606800
+- Zhao et al.
+(2025) Zixin Zhao, Damien Masson, Young-Ho Kim, Gerald Penn, and Fanny Chevalier.
+2025.
+Making the Write Connections: Linking Writing Support Tools with Writer’s Needs.
+In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (Yokohama, Kanagawa, Japan) (CHI ’25).
+Association for Computing Machinery, New York, NY, USA, Article 78, 21 pages.
+doi:10.1145/3706598.3713161
 - Zhu (2013) Xiaojin Zhu.
@@
 Persistent Homology: An Introduction and a New Text Representation for Natural Language Processing.
-In IJCAI International Joint Conference on Artificial Intelligence.
-1953–1959.
+In Proceedings of the Twenty-Third International Joint Conference on Artificial Intelligence.
+IJCAI, Beijing, China, 1953–1959.
 - Zhu and Kolassa (2018) Yaqian Zhu and John Kolassa.
@@
 2018), 2436–2453.
-https://doi.org/10.1080/03610918.2017.1348516
+doi:10.1080/03610918.2017.1348516
+Appendix A Appendix
+A.1. Extracting Information from the Text
+The visual representations are generated from information extracted from the text about the entities, locations, and events.
+The extraction follows three steps, in sequence: 1) extract the entities and their traits; 2) extract the locations; 3) extract the events for each sentence.
+For all steps, OpenAI’s “GPT4-o” model was used.
+A.1.1. Extracting Entities
+The prompt below extracts all the entities (i.e., characters and inanimate objects) from a story.
+<STORY TEXT>
+Extract all the entities in this story.
+For each entity, extract its ‘name’, an emoji best visually
+describing the entity (e.g., use the emoji of a person if it
+is a person but avoid reusing the same emojis),
+and properties about the entity, if any (no more than 3).
+Properties have to be adjectives describing the entity and
+their value should represent the intensity of the property
+(on a scale from 1 to 10).
+The model responds with a structured output using the following JSON schema:
+entities: [{ name: string, emoji: string properties: [{ name: string, value: number }] }]
+A.1.2. Extracting Locations
+The prompt below extracts all the locations from a story.
+<STORY TEXT>
+Extract all the main locations visited by the characters in this story. For each location, extract its ‘name’ and an emoji best visually representing the location
+The output is structured, using the following JSON schema:
+locations: [{ name: string, emoji: string }]
+A.1.3. Extracting Events
+Extracting the events is done after the entities and locations are extracted. First, the story is divided into sentences. Then, in parallel and for each sentence, the following prompt is used.
+BEFORE: <TEXT BEFORE>
+TEXT: <SENTENCE>
+Extract the actions done by the characters in TEXT and only the actions in TEXT.
+Do not extract the actions from BEFORE.
+Only consider actions that are happening exactly at the moment of TEXT, ignore memories etc.
+If there are no actions fulfilling these criteria in TEXT, then return an empty array.
+Source and target should be characters from this list: <ENTITIES>.
+Here are some possible locations but there might be others: <LOCATIONS> If an action is done by a character to itself, then the source and target character should be the same.
+For each action, extract the ‘name’ of the action (no more than 2 words), the source character (the character doing the action) and the target character (the character targeted by the action), and the location of the action (you can use ‘unknown’ if the location cannot be inferred from the text).
+With ¡TEXT BEFORE¿ corresponding to the text before the sentence (to give the model some context), ¡SENTENCE¿ the text of the sentence, ¡ENTITIES¿ the list of entities extracted previously, and ¡LOCATIONS¿ the list of locations extracted previously.
+The output is structured, using the following JSON schema:
+actions: [{ name: string, source: string, target: string, location: string }]
+All extracted events are associated with their respective sentences to support highlighting the text when manipulating the visual representations.
+Additionally, to make future extraction faster, only sentences that changed are re-extracted.
+A.2. Editing the Story by Manipulating the Visuals
+Below, we list the prompts used to suggest edits to the story when visual representations are modified.
+A.2.1. Reorder Events in the Timeline
+This prompt works by giving the list of events in the current and new order.
+Once the story is modified, the newly generated text is re-extracted fully to make sure the visual representations match the new text.
+<STORY TEXT>
+In this story, the current order of actions is:
+<CURRENT ORDER>
+Rewrite the story to EXACTLY follow this new order:
+<NEW ORDER>
+¡STORY¿ is the full text of the story, ¡CURRENT ORDER¿ is a list of the events in the story, as extracted previously, and ¡NEW ORDER¿ is this same list but with the events reorganized.
+Events are represented in the form “¡SOURCE¿ ¡EVENT NAME¿ ¡TARGET¿”.
+A.2.2. Adding, Changing or Removing An Action
+All the prompts resulting from manipulating links between nodes on the event view follow the same format:
+<STORY TEXT>
+SOURCE: <SOURCE ENTITY>
+TARGET: <TARGET ENTITY>
+ACTION: <ACTION NAME>
+<PROMPT>
+¡PROMPT¿ is either “Rewrite ¡blank¿ to add that SOURCE also ¡ACTION¿ TARGET” when adding an action, “Rewrite the story so that SOURCE also ¡ACTION¿ TARGET” when changing an action, and “Rewrite the story so that SOURCE does not do ACTION to TARGET” when removing an action.
+A.2.3. Removing an Entity
+<STORY TEXT>
+Rewrite the story so that there is no <ENTITY NAME>
+Note that no prompts are executed when adding an entity. This is only when adding an action that the text is modified.
+A.2.4. Moving an Entity
+<STORY TEXT>
+Rewrite the story so that <ENTITY NAME> never goes to the <CURRENT LOCATION> but instead goes to the <NEW LOCATION>
+A.2.5. Targeted Edits
+Visual edits can modify only a specific passage or set of scenes, if those scenes are selected on the timeline.
+For this to work, the prompts mentioned above are slightly modified so that ¡STORY TEXT¿ includes the fully story text with the selected passage replaced by “TEXT_TO_REWRITE”.
+Then, a new line is added defining “TEXT_TO_REWRITE” as the selected text passage.
+The prompt is then modified to indicate that it should only rewrite the passage.
+The result is then integrated into the full story.
+This approach is inspired by the prompts used in DirectGPT (Masson et al., 2024).